### v1

# Strong and Noise-Tolerant Entanglement in Dissipative Optomechanics

Abstract
Macroscopic entanglement, as critical quantum resources in quantum information science, has been extensively studied in optomechanical systems with purely dispersive coupling over the past decades. However, quantum entanglement, induced by purely dissipative coupling, remains unexplored. In this work, we study quantum entanglement in a Michelson–Sagnac interferometer, where the dispersive and the dissipative coupling can be arbitrarily switched on and off. With current experimental parameters, we demonstrate that the steady-state mechanical displacement exhibits a nonlinear (linear) dependence on the driving power with the purely dispersive (dissipative) coupling. Further, we find that quantum entanglement generated by purely dissipative coupling is significantly stronger and more noise-tolerant than that generated by purely dispersive coupling. When both couplings coexist, entanglement is weakened due to interference. Our findings provide a promising path to engineer strong and noise-tolerant quantum entanglement in purely dissipative quantum systems.
I Introduction
Due to its potential applications in gravitational wave detection Abbott et al. (2016) and quantum-classical boundary tests Haroche (1998), cavity optomechanics has garnered significant theoretical and experimental interest in quantum information science Aspelmeyer et al. (2014). With advances in ground-state cooling Liu et al. (2013); Teufel et al. (2011); Gröblacher et al. (2009); Schliesser et al. (2008), exploring quantum effects in cavity optomechanics has become a reality Lü et al. (2015); Safavi-Naeini et al. (2013); Purdy et al. (2013); Weis et al. (2010); Safavi-Naeini et al. (2011); Kronwald and Marquardt (2013); Koppenhöfer et al. (2023); Guo et al. (2016); Kyriienko et al. (2014a); Xiong et al. (2016, 2021a); Chen et al. (2021); Lü et al. (2013); Peng et al. (2023); Xiong et al. (2023); Wang et al. (2024); Mann et al. (2018); Jäger et al. (2019); Xiong et al. (2021b, 2022), especially for macroscopic quantum entanglement Vitali et al. (2007). Over the past two decades, numerous theoretical proposals have been put forward to achieve optomechanical entanglement Tian (2012, 2013); Wang and Clerk (2013); Chen et al. (2024, 2023); Liu et al. (2024a); Lai et al. (2022); Jiao et al. (2020, 2022); Liu et al. (2024b). Experimentally, macroscopic entanglement between two massive oscillators has also been realized in cavity optomechanics experimentally Ockeloen-Korppi et al. (2018). Notably, both theoretical and experimental proposals have predominantly relied on coherent dispersive coupling mechanisms between the cavity mode and the mechanical mode.
Very recently, unconventional dissipative optomechanics was first theoretically proposed in a superconducting microwave circuit Elste et al. (2009) and later extended to a Michelson-Sagnac interferometers (MSI) with a movable membrane Xuereb et al. (2011). In such dissipative optomechanics, the cavity linewidth depends on the mechanical displacement Elste et al. (2009), giving rise to the dissipative coupling between the mechanical and cavity modes. This unique coupling mechanism opens a potential avenue to investigate intriguing quantum phenomena, such as unconventional bistability Kyriienko et al. (2014b), quantum squeezing Qu and Agarwal (2015), enhanced ground-state cooling Huang and Chen (2018); Liu et al. (2023), quantum sensing Huang and Agarwal (2017), negative-damping instability Elste et al. (2009), self-sustained oscillations Huang et al. (2018), and phonon lasing Kuang et al. (2023); Zhang et al. (2022). Experimentally, dissipative optomechanics have been demonstrated in various platforms, including the MSI Sawadsky et al. (2015); Friedrich et al. (2011), microdisc resonator coupled to a nanomechanical waveguide Li et al. (2009), photonic crystal split-beam nanocavity Wu et al. (2014), graphene drums coupled to high-Q microsphere Cole et al. (2015), and levitated sphere trapped by dual-beam optical tweezers Kuang et al. (2023). It is evident that dissipative optomechanics has made significant progress both theoretically and experimentally. However, how dissipative optomechanics influences quantum entanglement, particularly under dissipative coupling or purely dissipative coupling conditions, remains an open and unexplored question.
In this work, we propose to investigate quantum entanglement in an experimentally feasible dissipative optomechanical system Sawadsky et al. (2015); Friedrich et al. (2011). The setup consists of a fixed perfect mirror and an effective movable mirror, realized by a Michelson–Sagnac interferometer (MSI) incorporating a membrane. Compared to dispersive optomechanics Vitali et al. (2007) with negligible dissipative coupling, the dispersive and dissipative couplings can be independently switched on and off by adjusting the membrane position or the reflectivity (transmissivity) of the beam splitter (BS) in this unconventional dissipative optomechanics Xuereb et al. (2011). We first examine the behavior of the steady-state mechanical displacement with purely dispersive or dissipative coupling. It is evident that that the displacement exhibits nonlinear dependence on the driving power with purely dispersive coupling, similar to conventional optomechanical systems Aspelmeyer et al. (2014). However, it becomes linear dependence of the driving power for purely dissipative coupling. We then study the effect of the purely dispersive or dissipative coupling on generation of quantum entanglement. We surprisely find that the quantum entanglement induced by purely dissipative coupling is not only much stronger than that generated by purely dispersive coupling, but also it is more robust aginst the bath temperature. For purely dispersive entanglement, its survival temperature is approximately $\sim 8.5\,\text{K}$, while for purely dissipative entanglement, it can be up to $\sim 25\,\text{K}$. When both couplings coexist, entanglement reduces due to quantum interference. Our work provides a promising path to generate strong and noise-resilient quantum entanglement in quantum systems with purely dissipative coupling.
This paper is organized as follows. In Sec. II, we introduce the proposed model and its Hamiltonian. Then, the steady-state solutions of the system are investigated in Sec. III. In Sec. IV, the dynamics of the fluctuation and covariance matrix are given. In Secs. V, quantum entanglement is investigated. Finally, a conclusion is given in Sec. VI.
II The model and Hamiltonian
Figure 1: (a) Schematic diagram of a typical dissipative optomechanical system, realized by a perfect fixed mirror $M_{1}$ and an effective movable mirror $\mathcal{M}$. (b) The structure of the effective movable mirror $\mathcal{M}$. The effective mirror consists of a Michelson-Sagnac interferometer (made up of a beam-splitter and two mirrors $M_{2}$ and $M_{3}$) and a movable membrane (mechanical resonator). $L$ is the effective length of the cavity ($2L$ is the length of the Sagnac mode ${\rm M_{1}-BS-M_{2}-M_{3}-BS}$). $a_{\rm in}$ is the input port and $a_{\rm out}$ is the output port.
As shown in Fig. 1(a), we consider an experimentally feasible dissipative optomechanics Sawadsky et al. (2015); Friedrich et al. (2011), consisting of a fixed mirror $M_{1}$ and an effective movable mirror $\mathcal{M}$. The effective mirror can be realized by a Michelson-Sagnac interferometer (MSI) and a movable membrane between mirrors $M_{2}$ and $M_{3}$ [see Fig. 1(b)]. $L$ is the effective length of the cavity ($2L$ is the length of the Sagnac mode ${\rm M}_{1}-{\rm BS}-{\rm M}_{2}-{\rm M}_{3}-{\rm BS}$). The membrane is treated quantum mechanically, with its displacement and momentum operators denoted by $x$ and $p$, respectively, satisfying the commutation relation $[x,p]=i\hbar$. Upon the introduction of an external driving field (with frequency $\omega_{d}$ and power $P$) at the beam splitter, the system can be effectively modeled as a typical dissipative optomechanical system. Thus, both the cavity frequency $\omega_{a}(x)$ and the decay rate $\kappa_{a}(x)$ depend on the displacement $x$ of the membrane. For small membrane displacements $x$, we can Taylor expand $\omega_{a}(x)$ and $\kappa_{a}(x)$ to the first order, i.e., $\omega(x)\approx\omega_{a}-g_{\omega}x$ and $\kappa(x)\approx\kappa_{a}-g_{\kappa}x$, where $\omega_{a}$ ($\kappa_{a}$) is the frequency (decay rate) of the cavity with annihilation (creation) operator $a$ ($a^{\dagger}$) when the membrane is at its equilibrium. The Hamiltonian of the considered system is given by Elste et al. (2009); Xuereb et al. (2011)
$H_{S}=$ $(\omega_{a}-g_{\omega}x)a^{\dagger}a+\frac{1}{2}\omega_{m}(x^{2}+p^{2})$
[…]
where $\omega_{m}$ is the frequency of the membrane, $\omega$ is the frequency of the bosonic mode (i.e., bath) with annihilation (creation) operator $a_{\omega}$ ($a_{\omega}^{\dagger}$), $x^{\prime}=\sqrt{2\kappa_{a}}-{g_{\kappa}x}/{\sqrt{2\kappa_{a}}}$ is the position-dependent coupling strength between the optical cavity and the bosonic mode. In fact, the bosonic mode here is to introduce a zero-value expectation input field $a_{\rm in}=\int\frac{d\omega}{\sqrt{2\pi}}e^{-i\omega(t-t_{0})}a_{\omega,0}$, with $a_{\omega,0}$ the operator at an initial time, to the cavity mode $a$, which can be easily obtained following the standard input-output theory with the Markov approximation. At the optical domain, the input field satisfies the relations walls1994gj
$\langle a_{\rm in}(t)a_{\rm in}(t^{\prime})\rangle=$ $0,$
$\langle a_{\rm in}(t)a^{\dagger}_{\rm in}(t^{\prime})\rangle=$ $\delta(t-t^{\prime}).$ (2)
Substituting the input field $a_{\rm in}$ into Eq. (II), the Hamiltonian of the driven unconventional MSI, with respect to the driving frequency $\omega_{d}$, becomes
$\mathcal{H}_{S}=$ $(\Delta_{a}-g_{\omega}x)a^{\dagger}a+\frac{1}{2}\omega_{m}(x^{2}+p^{2})$
$+ix^{\prime}(A_{\rm in}a^{\dagger}-A_{\rm in}^{\dagger}a),$ (3)
where $\Delta_{a}=\omega_{a}-\omega_{d}$ is the frequency detuning of the cavity mode from the driving field, and
$A_{\rm in}=a_{\rm in}+\mathcal{E}.$ (4)
Here $\mathcal{E}=|\mathcal{E}|\exp(i\theta)$, with $|\mathcal{E}|=\sqrt{\mathcal{P}/\hbar\omega_{d}}$ being the amplitude of the driving field and $\theta$ the phase. Notably, $\mathcal{E}$ can also be interpreted as the steady-state value of the operator $A_{\rm in}$, while $a_{\rm in}$ represents the fluctuation of $A_{\rm in}$.
III Steady-state solution
When dissipation is included, the dynamics of the proposed system can be governed by the quantum Langevin equantion $\mathcal{O}=-i[\mathcal{O},\mathcal{H}_{S}]+{Dissipation}+{Noise}$, specifically,
$\dot{x}=$ $\omega_{m}p,$
$\dot{p}=$ $g_{\omega}a^{\dagger}a-\omega_{m}x+\frac{ig_{\kappa}}{\sqrt{2\kappa_{a}}}(A_{\rm in}^{\dagger}a-A_{\rm in}a^{\dagger})-\gamma_{m}p+\xi,$
$\dot{a}=$ $-(\kappa_{a}+i\Delta_{a})a+ig_{\omega}ax+g_{\kappa}ax+x^{\prime}A_{\rm in},$ (5)
where $\gamma_{m}$ is the damping rate of the membrane and $\xi$ is the Brownian noise operator arising from the coupling of the membrane oscillator to the hot environment. The correlation function of $\xi$ is given by Giovannetti and Vitali (2001) $\langle\xi(t)\xi(t^{\prime})\rangle=\frac{\gamma_{m}}{\omega_{m}}\int\frac{d\omega}{2\pi}e^{-i\omega(t-t^{\prime})}\omega[{\rm coth}(\frac{\hbar\omega}{2k_{b}T})+1]$. In the limit of the high quality factor, $\omega_{m}\gg\gamma_{m}$, the Brownian noise $\xi$ becomes am approximate Markovian noise Benguria and Kac (1981), i.e.,
$\langle\xi(t)\xi(t^{\prime})+\xi(t^{\prime})\xi(t)\rangle/2\backsimeq\gamma_{m}(2n_{\rm th}+1)\delta(t-t^{\prime})$ (6)
where the mean thermal phonon number $n_{\rm th}$ = $[{\rm exp}(\hbar\omega_{m}/{k_{B}T}-1)]^{-1}$ with $k_{B}$ the Boltzmann constant and $T$ the bath temperature.
Figure 2: The normalized steady-state displacment $x_{s}$ of the membrane vs (a) the intracavity mean photon number $n_{s}=|a_{s}|^{2}$ and (b) the power of the driving field. The red curve corresponds to the case of the purely dispersive coupling ($g_{\omega}/2\pi=2$ Hz). The blue curve is plotted for the purely dissipative coupling ($g_{\kappa}/2\pi=2$ Hz). In (a) and (b), $\Delta_{a}=3\kappa_{a}$ is assumed.
When the evolution time is long enough, the system reaches its steady state, denoted by $a_{s}$, $x_{s}$, and $p_{s}$. From Eq. (III), the steady-state equation can be directly obtained as
$\dot{x}_{s}=$ $\omega_{m}p_{s},$
$\dot{p}_{s}=$ $g_{\omega}|a_{s}|^{2}-\omega_{m}x_{s}+\frac{ig_{\kappa}}{\sqrt{2\kappa_{a}}}(\mathcal{E}^{*}a_{s}-\mathcal{E}a_{s}^{*})-\gamma_{m}p_{s},$
$\dot{a}_{s}=$ $-(\kappa_{s}+i{\Delta}_{s})a_{s}+x_{s}^{\prime}\mathcal{E},$ (7)
where ${\Delta}_{s}=\Delta_{a}-{g_{\omega}}x_{s}$, $\kappa_{s}=\kappa_{a}-g_{\kappa}x_{s}$, and $x_{s}^{\prime}=\sqrt{2\kappa_{a}}-{g_{\kappa}x_{s}}/{\sqrt{2\kappa_{a}}}$. At the steady state, we have $\dot{\mathcal{O}}_{s}=0$, equivalently
$\omega_{m}p_{s}=$ $0,$
$g_{\omega}|a_{s}|^{2}-\omega_{m}x_{s}+\frac{ig_{\kappa}}{\sqrt{2\kappa_{a}}}(\mathcal{E}^{*}a_{s}-\mathcal{E}a_{s}^{*})-\gamma_{m}p_{s}=$ $0,$ (8)
$-(\kappa_{s}+i{\Delta}_{s})a_{s}+x_{s}^{\prime}\mathcal{E}=$ $0$
By solving these equations, we have
$p_{s}=$ $0,~{}~{}a_{s}=\frac{x_{s}^{\prime}\mathcal{E}}{\kappa_{s}+i{\Delta}_{s}},$
$x_{s}=$ $\frac{g_{\omega}n_{s}}{\omega_{m}}+\frac{ig_{\kappa}}{\sqrt{2\kappa_{a}}}\cdot\frac{a_{s}\mathcal{E}^{*}-a_{s}^{*}\mathcal{E}}{\omega_{m}}.$ (9)
Obivously, the steady-state displacement of the membrane is determined by both the dispersive and dissipative coupling between photons and phonons. For a purely dispersive coupling ($g_{\kappa}=0$), $x_{s}$ is proportional to the intracavity mean photon number $n_{s}=|a_{s}|^{2}$ [see the red curve in Fig. 2(a)]. For a purely dissipative coupling ($g_{\omega}=0$), $x_{s}$ becomes
$x_{s}=[1-\sqrt{1-2(g_{\kappa}/\kappa_{a})^{2}(\Delta_{a}/\omega_{m})n_{s}}]\omega_{m}\kappa_{a},$ (10)
where $x_{s}\neq 2\kappa_{a}/g_{\kappa}$ and $n_{s}\leq(\kappa_{a}/g_{\kappa})^{2}(\omega_{m}/2\Delta_{a})$ are both required. When $n_{s}=0$, $x_{s}=0$, i.e., the membrane is positioned at its equilibrum. Here $x_{s}$ and $n_{s}$ are quadratic dependence for the purely dissipative coupling, demonstrated by the blue curve in Fig. 2(a). One may further notice that the displacement of the membrane caused by the purely dispersive coupling is larger than that caused by the purely dissipative coupling, as clearly shown in Fig. 2(a). To plot Fig. 2(a), we adopt typical experimental parameters Sawadsky et al. (2015): the wavelength of the driving field $\lambda=1064$ nm, $\omega_{d}/2\pi=c/\lambda=281.96$ THz, the quality factor of the membrane $\mathcal{Q}_{m}=5.8\times 10^{5}$, $\omega_{m}/2\pi=136$ kHz, $\gamma_{m}/2\pi=0.23$ Hz, $L=8.7$ cm, $m=80$ ng, and $\kappa_{a}/2\pi=1.5$ MHz. These experimental parameters ensure the stability of the proposed system according to Routh-Hurwitz criterion.
From Eq. (III), one can see that $x_{s}$ actually satisfies the following cubic equation related to the amplitude of the driving field $\mathcal{E}$,
$ax_{s}^{3}+bx_{s}^{2}+cx_{s}+d=0,$ (11)
where
$a=$ $2(g_{\omega}^{2}+g_{\kappa}^{2})\omega_{m}\kappa_{a},$
$b=$ $-4\omega_{m}\kappa_{a}(g_{\omega}\Delta_{a}+g_{\kappa}\kappa_{a})-3g_{\omega}g_{\kappa}^{2}|\mathcal{E}|^{2},$
$c=$ $2\omega_{m}\kappa_{a}(\kappa_{a}^{2}+\Delta_{a}^{2})+2g_{\kappa}|\mathcal{E}|^{2}(g_{\kappa}\Delta_{a}+4g_{\omega}\kappa_{a}),$ (12)
$d=$ $-4\kappa_{a}|\mathcal{E}|^{2}(g_{\omega}\kappa_{a}+g_{\kappa}\Delta_{a}).$
Explicitly, Eq. (11) may have three real roots, corresponding to bistability. But in fact, the bistability can only be predicted for puly dispersive coupling, as shown by the red curve in Fig. 2(b). This indicates that the purely dissipative coupling can not generate the typical stability in optomechanics, demonstrated by the blue curve in Fig. 2(b).
IV Covariance matrix
With steady-state values in hand, we further expand each operator $\mathcal{O}$ around its steady state, i.e., $\mathcal{O}=\delta\mathcal{O}+\mathcal{O}_{s}$. Substituting this expression back into Eq. (III), we then obtain a set of equations for fluctuations ($\delta\mathcal{O}$). For the strong driving field ($|a_{s}|^{2}\gg 1$), the higher-order fluctuation terms can be safely neglected Vitali et al. (2007), corresponding to linearization. As a result, the dynamics for fluctuations can be written as
$\dot{\delta x}=$ $\omega_{m}\delta p,$
$\dot{\delta p}=$ $\frac{G_{\omega}\delta a^{\dagger}+G_{\omega}^{*}\delta a}{\sqrt{2}}-\omega_{m}\delta x-\gamma_{m}\delta p+\xi$
$+\frac{ig_{\kappa}(\mathcal{E}^{*}\delta a-\mathcal{E}\delta a^{\dagger})}{\sqrt{2\kappa_{a}}}+\frac{i(G_{\kappa}\delta a_{\rm in}^{\dagger}-G_{\kappa}^{*}\delta a_{\rm in})}{2\sqrt{\kappa_{a}}},$ (13)
$\dot{\delta a}=$ $-(\kappa_{s}+i\Delta_{s})\delta a+(\frac{G_{\kappa}+iG_{\omega}}{\sqrt{2}}-\frac{g_{\kappa}\mathcal{E}}{\sqrt{2\kappa_{a}}})\delta x+x_{s}^{\prime}a_{\rm in},$
where $G_{\omega(\kappa)}=\sqrt{2}g_{\omega(\kappa)}a_{s}$ is the linearized dispersive (dissipative) coupling strength, enhanced by the amplitude of the driving field. By further defining
$\delta x_{a}=$ $\frac{\delta a^{\dagger}+\delta a}{\sqrt{2}},~{}~{}\delta p_{a}=i\frac{\delta a^{\dagger}-\delta a}{\sqrt{2}},$
$x_{\rm in}=$ $\frac{a_{\rm in}^{\dagger}+a_{\rm in}}{\sqrt{2}},~{}~{}p_{\rm in}=i\frac{a_{\rm in}^{\dagger}-a_{\rm in}}{\sqrt{2}},$ (14)
Eq. (IV) can be rewritten in a more compact form, i.e.,
$\dot{u}(t)=Au(t)+n(t),$ (15)
where $u^{T}(t)=(\delta x,\delta p,\delta x_{a},\delta p_{a})$ is the vector of fluctuation operators, $n(t)$ is the vector of noise operator, given by $n^{T}(t)=(0,\xi+\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}p_{\rm in},x_{s}^{\prime}x_{\rm in},x_{s}^{\prime}p_{\rm in})$, and the drift matrix $A$ is
$A=\left(\begin{array}[]{cccc}0&\omega_{m}&0&0\\ -\omega_{m}&-\gamma_{m}&G_{\omega}+\frac{G_{\kappa}\Delta_{s}}{\sqrt{2\kappa_{a}}x_{s}^{\prime}}&-\frac{G_{\kappa}\kappa_{s}}{\sqrt{2\kappa_{a}}x_{s}^{\prime}}\\ G_{\kappa}-\frac{2G_{\kappa}\kappa_{s}}{\sqrt{2\kappa_{a}}x_{s}^{\prime}}&0&-\kappa_{s}&\Delta_{s}\\ G_{\omega}-\frac{2G_{\kappa}\Delta_{s}}{\sqrt{2\kappa_{a}}x_{s}^{\prime}}&0&-\Delta_{s}&-\kappa_{s}\\ \end{array}\right).$ (16)
To obtain the drift matrix $A$, we have taken $a_{s}$ as real number, which can be realized by setting $\tan\theta=\Delta_{a}/\kappa_{a}$ in Eq. (III).
Since the phonon noise $\xi$ and the photon noise $a_{\rm in}$ are zero-mean quantum Gaussian noises and the dynamics is linearized [see Eq. (IV)], the quantum steady state for fluctuations is a zero-mean bipartite Gaussian state, fully described by a $4\times 4$ correlation matrix ${V}$, where $V_{ij}=\langle u_{i}(t)u_{j}(t^{\prime})+u_{j}(t^{\prime})u_{i}(t)\rangle/2$ with $i,j=1,2,3,4$. The $V$ is obtained by solving the Lyapunov equation $AV+VA^{T}=-D$, where $D$ is the diffusion matrix, defined by $\langle n_{i}(t)n_{j}(t^{\prime})+n_{j}(t^{\prime})n_{i}(t)\rangle/2=D_{ij}\delta(t-t^{\prime})$. More specifically, $D$ can be expressed as
$D=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&\gamma_{m}(2n_{\rm th}+1)+\frac{G_{\kappa}^{2}}{4\kappa_{a}}&0&\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}\frac{x_{s}^{\prime}}{2}\\ 0&0&\frac{x_{s}^{\prime,2}}{2}&0\\ 0&\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}\frac{x_{s}^{\prime}}{2}&0&\frac{x_{s}^{\prime,2}}{2}\\ \end{array}\right).$ (17)
Compared to the case of purely dispersive coupling, the dissipative coupling strength appears in the off-diagonal elements of the diffusion matrix, rendering $D$ non-diagonal.
To measure optomechanical entanglement, the logarithmic negativity ($E_{N}$) is taken Adesso et al. (2004), i.e.,
$E_{N}\equiv\rm max[0,-ln2\eta^{-}],$ (18)
where $\eta^{-}=2^{-1/2}[\Sigma({V})-(\Sigma({V})^{2}-4\rm det{{V}})^{1/2}]^{1/2}$ with $\Sigma({V})=\rm{det}{A}+\rm{det}B-2\rm{det}C$, ${{V}}=\begin{pmatrix}A&C\\ {C^{T}}&B\\ \end{pmatrix}$ is the $2\times 2$ block form. When $E_{N}>0~{}(=0)$, the state is entangled (separable). Moreover, the larger $E_{N}$, the stronger entanglement.
V Optomechanical Entanglement
To investigate optomechanical entanglement, we consider three specific scenarios: (i) purely dispersive coupling, (ii) purely dissipative coupling, and (iii) the coexistence of both couplings.
V.1 Purely dispersive coupling ($g_{\kappa}=0$)
In the purely dispersive coupling regime ($g_{\kappa}=0$), we present the entanglement measure $E_{N}$ as a function of both the dispersive coupling strength $g_{\omega}$ and the normalized effective cavity detuning $\Delta_{s}/\omega_{m}$ in Fig. 3(a) to explore the conditions for optimal optomechanical entanglement at a bath temperature of $T=0.4$ K. The results reveal that the generated entanglement remains relatively weak compared to that typically observed in conventional optomechanical systems. This is primarily due to the inherent limitation in the dispersive coupling strength $g_{\omega}$ (or equivalently $G_{\omega}$), which must remain small within the stable regime to avoid system instability, thereby restricting the achievable entanglement. Nonetheless, optimal entanglement can still be achieved around the point ($\Delta_{s}/\omega_{m},g_{\omega}/2\pi)\approx(2,3.1)$, as marked by the green star, indicating that even under these constraints, entanglement can be optimized.
Figure 3: (a) The logarithmic negativity $E_{N}$ vs the normalized detuning $\Delta/\omega_{m}$ and the purely dispersive coupling strength $g_{\omega}$ at bath temperature $T=400$ mK. (b) The logarithmic negativity $E_{N}$ vs the dispersive coupling strength $g_{\omega}$ with $T=400$ mK (see the red curve) and $T=4$ K (see the blue dashed curve), where $\Delta_{s}/\omega_{m}=2$ is fixed. (c) The logarithmic negativity $E_{N}$ vs the bath temperature $T$ with $g_{\omega}/2\pi=3.1$ Hz and $\Delta_{s}/\omega_{m}=2$. Other parameters are the same as those in Fig. 2 except for $\mathcal{P}=50$ mW.
To further investigate the role of dispersive coupling strength on entanglement, we plot $E_{N}$ as a function of $g_{\omega}$ in Fig. 3(b), where $\Delta_{s}\approx 2\omega_{m}$ is fixed. It is obviously shown that $E_{N}$ increases with $g_{\omega}$. Once up to its maximum, $E_{N}$ will have a sharp reduction by further increasing $g_{\omega}$. Such abrupt behavior corresponds to the transition from the stability to instability, as shown by the pink-shaded region. Additionally, when the bath temperature increases to $T=4$ K [see the blue dashed curve], the coupling for generation of quantum entanglement is required to be larger and $E_{N}$ has a significant reduction. This indicates that quantum entanglement can be affected by the bath temperature. To clearly evaluate the effect of bath temperature, we plot $E_{N}$ as a function of the bath temperature in Fig. 3(c), where $g_{\omega}/2\pi=3.1$ Hz is fixed to obtain optimal entanglement. As the temperature increases, $E_{N}$ gradually declines. When the bath temperature increases to $T\approx 8.5$ K, $E_{N}=0$, that is, quantum entanglment vanishes and the system becomes separable.
V.2 Purely dissipative coupling ($g_{\omega}=0$)
Figure 4: (a) The logarithmic negativity $E_{N}$ vs the normalized detuning $\Delta/\omega_{m}$ and the purely dissipative coupling strength $g_{\kappa}$ at bath temperature $T=400$ mK. (b) The logarithmic negativity $E_{N}$ vs the dissipative coupling strength $g_{\kappa}$ with $T=400$ mK (see the red curve) and $T=4$ K (see the blue dashed curve), where $\Delta_{s}/\omega_{m}=0.1$ is fixed. (c) The logarithmic negativity $E_{N}$ vs the bath temperature $T$ with $g_{\kappa}/2\pi=19$ Hz and $\Delta_{s}/\omega_{m}=0.1$. Other parameters are the same as those in Fig. 2 except for $\mathcal{P}=50$ mW.
In the purely dissipative case ($g_{\omega}=0$), we observe that the optimal entanglement can be significantly stronger than in the purely dispersive case, even surpassing the entanglement achieved in conventional optomechanical systems (i.e., $E_{N}=0.35$). This is demonstrated in Fig. 4(a), where $E_{N}$ is plotted as a function of both the dissipative coupling strength ($g_{\kappa}$) and the normalized effective cavity detuning $\Delta_{s}/\omega_{m}$. The optimal entanglement is found at approximately ($\Delta_{s}/\omega_{m},g_{\kappa}/2\pi)\approx(0.1,20)$, as indicated by the green star. In contrast to the purely dispersive case [see Fig. 2(a)], where the optimal entanglement occurs near the mechanical red sideband at $\Delta_{s}/\omega_{m}=2$, the optimal entanglement in the purely dissipative case is obtained when the cavity is nearly resonant with the driving field, specifically at $\Delta_{s}/\omega_{m}=0.1$, which is far removed from the mechanical red sideband.
Furthermore, in Fig. 4(b), we investigate the behavior of the optomechanical entanglement as a function of the dissipative coupling strength, while fixing $\Delta_{s}=0.1\omega_{m}$. The resulting behavior of $E_{N}$ mirrors that observed in the purely dispersive coupling case (see the red curve). The abrupt drop in entanglement marks the transition of the system from a stable to an unstable region, as indicated by the pink area. Notably, a relatively large dissipative coupling strength is required to achieve the optimal entanglement. This suggests that in the regime dominated by dissipative coupling, stronger interactions are necessary to generate significant quantum effects. As the bath temperature increases, for instance to $T=4$ K, the entanglement is reduced (see the blue curve). To explore the influence of bath temperature on optomechanical entanglement, we fix $g_{\kappa}/2\pi=19$ Hz and present the results in Fig. 4(c). We observe a sharp decrease in entanglement as the system temperature increases from 0 K to 5 K. However, beyond this point, the entanglement becomes more tolerant to further temperature increases, suggesting that the dissipative coupling becomes more robust to thermal noise at higher temperatures. Remarkably, entanglement persists even up to 25 K, This indicates that strong, noise-resistant entanglement can be generated through purely dissipative coupling, opening new possibilities for quantum information processing and communication in noisy
Figure 5: (a) The logarithmic negativity $E_{N}$ vs the normalized detuning $\Delta/\omega_{m}$ and the ratio of $g_{\kappa}/g_{\omega}$ at bath temperature $T=400$ mK. (b) The logarithmic negativity $E_{N}$ vs the ratio of $g_{\kappa}/g_{\omega}$ with $T=400$ mK (see the red curve) and $T=4$ K (see the blue dashed curve), where $\Delta_{s}/\omega_{m}=2.2$ is fixed. (c) The logarithmic negativity $E_{N}$ vs the bath temperature $T$ with $g_{\kappa}/g_{\omega}=0.01$ (see the red curve) and $g_{\kappa}/g_{\omega}=0.1$ K (see the blue dashed curve), where $\Delta_{s}/\omega_{m}=2.2$. Other parameters are the same as those in Fig. 2 except for $\mathcal{P}=50$ mW, $g_{\omega}/2\pi=3$ Hz.
V.3 Combined coupling ($g_{\omega}\neq 0$ and $g_{\kappa}\neq 0$)
To examine the interplay between dispersive and dissipative coupling effects on optomechanical entanglement, we set the dispersive coupling strength to $g_{\omega}/2\pi=3$ Hz, corresponding to the optimal entanglement condition at $\Delta_{s}/\omega_{m}=2$ in the purely dispersive coupling regime [refer to Fig. 3(a)]. By progressively increasing the dissipative coupling strength, we observe a gradual decline in the entanglement, as shown in Fig. 5(a). The reduction in entanglement possibly arises from the interference between the dispersive and dissipative couplings. When the bath temperature increases to $T=4$ K, it is observed that the optomechancial entanglement is further suppressed, as depicted by the blue dashed curve in Fig. 5(b). Moreover, the larger $g_{\kappa}$ is not benificial to entanglement generation. This can also be revealed by Fig. 5(c), where $g_{\kappa}/g_{\omega}=0.01$ (the red curve) and $g_{\kappa}/g_{\omega}=0.1$ (the blue curve) are taken . For $g_{\kappa}/g_{\omega}=0.1$, the survival temperature of the entanglement is about $5$ K. But for $g_{\kappa}/g_{\omega}=0.01$, the survival temperature can be up to $8$ K.
VI Conclusion
In this paper, we investigate quantum entanglement in a dissipative optomechanical system, acted by a fixed mirror and a Michelson–Sagnac interferometer with a membrane. This system allows the dissipative and dispersive couplings to be independently tuned on and off by adjusting the membrane position or the reflectivity (transmissivity) of the beam splitter, enabling us to explore quantum entanglement in three distinct scenarios: purely dispersive case, purely dissipative case, and combined coupling case. Our analysis reveals that, in the purely dispersive case, the steady-state mechanical displacement exhibits a nonlinear dependence on the driving power, whereas in the purely dissipative case, the displacement shows a linear dependence. Furthermore, entanglement generated through purely dissipative coupling is not only significantly stronger but also more noise-tolerant compared to that achieved via purely dispersive coupling. When both coupling mechanisms coexist, quantum entanglement is reduced, likely due to the interference between the two coupling interactions. These findings suggest the potential of purely dissipative coupling to engineer strong and noise-tolerant quantum entanglement, offering a promising pathway for advancing quantum technologies.
This work is supported by the National Natural Science Foundation of China (Grant No. 12175001 ), the “Pioneer” and “Leading Goose” R&\&&D Program of Zhejiang (Grant No. 2025C01028), the Natural Science Foundation of Zhejiang Province (Grant No. LY24A040004), and the Shenzhen International Quantum Academy (Grant No. SIQA202412).

### v2

@@
 However, quantum entanglement, induced by purely dissipative coupling, remains unexplored.
-In this work, we study quantum entanglement in a Michelson–Sagnac interferometer, where the dispersive and the dissipative coupling can be arbitrarily switched on and off.
+In this work, we study quantum entanglement in a Michelson–Sagnac interferometer, where the dipersive and the dissipative coupling can be arbitrarily switched on and off.
 With current experimental parameters, we demonstrate that the steady-state mechanical displacement exhibits a nonlinear (linear) dependence on the driving power with the purely dispersive (dissipative) coupling.
@@
 (2020, 2022); Liu et al.
-(2024b).
+(2024b); Shang and Li (2024).
 Experimentally, macroscopic entanglement between two massive oscillators has also been realized in cavity optomechanics experimentally Ockeloen-Korppi et al.
@@
 Our work provides a promising path to generate strong and noise-resilient quantum entanglement in quantum systems with purely dissipative coupling.
-This paper is organized as follows.
+The rest paper is organized as follows.
 In Sec.
@@
 In fact, the bosonic mode here is to introduce a zero-value expectation input field $a_{\rm in}=\int\frac{d\omega}{\sqrt{2\pi}}e^{-i\omega(t-t_{0})}a_{\omega,0}$, with $a_{\omega,0}$ the operator at an initial time, to the cavity mode $a$, which can be easily obtained following the standard input-output theory with the Markov approximation.
-At the optical domain, the input field satisfies the relations walls1994gj
+At the optical domain, the input field satisfies the relations Walls and Milburn (1994)
 $\langle a_{\rm in}(t)a_{\rm in}(t^{\prime})\rangle=$ $0,$
@@
 In this paper, we investigate quantum entanglement in a dissipative optomechanical system, acted by a fixed mirror and a Michelson–Sagnac interferometer with a membrane.
-This system allows the dissipative and dispersive couplings to be independently tuned on and off by adjusting the membrane position or the reflectivity (transmissivity) of the beam splitter, enabling us to explore quantum entanglement in three distinct scenarios: purely dispersive case, purely dissipative case, and combined coupling case.
+This system allows the dissipative and dispersive couplings to be independently tuned on and off by adjusting the membrane position or the reflectivity (transmissivity) of the BS, enabling us to explore quantum entanglement in three distinct scenarios: purely dispersive case, purely dissipative case, and combined coupling case.
 Our analysis reveals that, in the purely dispersive case, the steady-state mechanical displacement exhibits a nonlinear dependence on the driving power, whereas in the purely dissipative case, the displacement shows a linear dependence.
@@
 LY24A040004), and the Shenzhen International Quantum Academy (Grant No.
-SIQA202412).
+SIQA2024KFKT010).

### v3

@@
 Abstract
-Macroscopic entanglement, as critical quantum resources in quantum information science, has been extensively studied in optomechanical systems with purely dispersive coupling over the past decades.
-However, quantum entanglement, induced by purely dissipative coupling, remains unexplored.
-In this work, we study quantum entanglement in a Michelson–Sagnac interferometer, where the dipersive and the dissipative coupling can be arbitrarily switched on and off.
-With current experimental parameters, we demonstrate that the steady-state mechanical displacement exhibits a nonlinear (linear) dependence on the driving power with the purely dispersive (dissipative) coupling.
-Further, we find that quantum entanglement generated by purely dissipative coupling is significantly stronger and more noise-tolerant than that generated by purely dispersive coupling.
-When both couplings coexist, entanglement is weakened due to interference.
-Our findings provide a promising path to engineer strong and noise-tolerant quantum entanglement in purely dissipative quantum systems.
+Macroscopic entanglement, as critical quantum resources in quantum information science, has been extensively studied in coherent optomechanics over the past decades.
+However, entanglement in dissipative optomechanics, where the cavity linewidth is dependent on the position of the mechanical resonator, remains unexplored.
+In this work, we study quantum entanglement in a dissipative optomechanics, realized by a Michelson–Sagnac interferometer and a movable membrane.
+Such dissipative optomechanics allows to switch coherent and dissipative optomechanical coupling at will.
+With reliable parameters, we demonstrate that the steady-state mechanical displacement exhibits a nonlinear (linear) dependence on the driving power with coherent (dissipative) coupling.
+Further, we show that quantum entanglement generated by dissipative coupling is significantly stronger and more noise-tolerant than that generated by coherent coupling.
+When both the coherent and dissipative couplings are simultaneously considered, entanglement is instead weakened due to quantum interference.
+Our result indicates that dissipative optomechanical coupling may be a promising path to engineer strong and noise-tolerant quantum entanglement.
 I Introduction
-Due to its potential applications in gravitational wave detection Abbott et al.
-(2016) and quantum-classical boundary tests Haroche (1998), cavity optomechanics has garnered significant theoretical and experimental interest in quantum information science Aspelmeyer et al.
+Coherent optomechanics, where the cavity frequency is dependent on the position of the mechanical resonator, has garnered siginificant interest theoretically and experimentally over the past two decades Aspelmeyer et al.
 (2014).
-With advances in ground-state cooling Liu et al.
-(2013); Teufel et al.
+In such systems, various proposals have been put forward, including ground-state cooling Teufel et al.
 (2011); Gröblacher et al.
 (2009); Schliesser et al.
-(2008), exploring quantum effects in cavity optomechanics has become a reality Lü et al.
+(2008), bistability Dorsel et al.
+(1983); Gozzini et al.
+(1985) and tristability Xiong et al.
+(2016), optomechancally induced transparency Weis et al.
 (2010); Safavi-Naeini et al.
+(2011); Kronwald and Marquardt (2013), quantum squeezing Liao and Law (2011); Nunnenkamp et al.
+(2010); Lü et al.
+(2015); Aggarwal et al.
 (2020); Purdy et al.
-(2013); Weis et al.
-(2010); Safavi-Naeini et al.
-(2011); Kronwald and Marquardt (2013); Koppenhöfer et al.
-(2023); Guo et al.
-(2016); Kyriienko et al.
-(2014a); Xiong et al.
-(2016, 2021a); Chen et al.
-(2021); Lü et al.
+(2013); Luo et al.
+(2022), phase transition Wang et al.
+(2024); Jäger et al.
+(2019); Mann et al.
+(2018); Xu et al.
+(2017), microwave-optical conversion Guo et al.
+(2016); Han et al.
+(2021); Sahu et al.
+(2022); Andrews et al.
+(2014); Chen et al.
+(2023a); Xu et al.
+(2016), spin-optomechanical interface Xiong et al.
+(2021a); Chen et al.
 (2021); Peng et al.
 (2023); Xiong et al.
-(2023); Wang et al.
-(2024); Mann et al.
 (2023); Jäger et al.
-(2019); Xiong et al.
-(2021b, 2022), especially for macroscopic quantum entanglement Vitali et al.
-(2007).
-Over the past two decades, numerous theoretical proposals have been put forward to achieve optomechanical entanglement Tian (2012, 2013); Wang and Clerk (2013); Chen et al.
-(2024, 2023); Liu et al.
+(2019), higher-order exceptional point Xiong et al.
+(2021b, 2022), and macroscopic quantum entanglement Vitali et al.
+(2007); Tian (2012, 2013); Wang and Clerk (2013); Chen et al.
+(2024, 2023b); Liu et al.
 (2024a); Lai et al.
@@
 (2020, 2022); Liu et al.
-(2024b); Shang and Li (2024).
-Experimentally, macroscopic entanglement between two massive oscillators has also been realized in cavity optomechanics experimentally Ockeloen-Korppi et al.
+(2024b); Shang and Li (2024); Ockeloen-Korppi et al.
 (2018).
-Notably, both theoretical and experimental proposals have predominantly relied on coherent dispersive coupling mechanisms between the cavity mode and the mechanical mode.
-Very recently, unconventional dissipative optomechanics was first theoretically proposed in a superconducting microwave circuit Elste et al.
-(2009) and later extended to a Michelson-Sagnac interferometers (MSI) with a movable membrane Xuereb et al.
-(2011).
-In such dissipative optomechanics, the cavity linewidth depends on the mechanical displacement Elste et al.
-(2009), giving rise to the dissipative coupling between the mechanical and cavity modes.
-This unique coupling mechanism opens a potential avenue to investigate intriguing quantum phenomena, such as unconventional bistability Kyriienko et al.
-(2014b), quantum squeezing Qu and Agarwal (2015), enhanced ground-state cooling Huang and Chen (2018); Liu et al.
-(2023), quantum sensing Huang and Agarwal (2017), negative-damping instability Elste et al.
-(2009), self-sustained oscillations Huang et al.
-(2018), and phonon lasing Kuang et al.
-(2023); Zhang et al.
-(2022).
-Experimentally, dissipative optomechanics have been demonstrated in various platforms, including the MSI Sawadsky et al.
+Recently, dissipative optomechanics was theoretically proposed in a superconducting microwave circuit Elste et al.
+(2009), where the cavity linewidth is a function of the mechanical displacement Elste et al.
+(2009), giving rise to a dissipative coupling between the mechanical and cavity modes.
+Later, the setup was extended to a Michelson-Sagnac interferometers (MSI) with a movable membrane Xuereb et al.
+(2011), where coherent or dissipative optomechanical coupling is allowed to switch on and off at will by tuning the position of the membrane or the reflectivity of the beam splitter.
+Experimentally, dissipative optomechanics have been demonstrated in MSI Sawadsky et al.
 (2015); Friedrich et al.
@@
 (2023).
-It is evident that dissipative optomechanics has made significant progress both theoretically and experimentally.
-However, how dissipative optomechanics influences quantum entanglement, particularly under dissipative coupling or purely dissipative coupling conditions, remains an open and unexplored question.
+Within dissipative optomechanics, diverse intriguing quantum phenomena, such as unconventional bistability Kyriienko et al.
+(2014), quantum squeezing Qu and Agarwal (2015); Huang and Chen (2020), enhanced ground-state cooling Huang and Chen (2018); Liu et al.
+(2023), quantum sensing Huang and Agarwal (2017), negative-damping instability Elste et al.
+(2009), self-sustained oscillations Huang et al.
+(2018), and phonon lasing Kuang et al.
+(2023); Zhang et al.
+(2022) have been investigated.
+However, macroscopic quantum entanglement has not been revealed yet with dissipative optomechanical coupling, although it has been intensively studied in coherent optomechanics.
 In this work, we propose to investigate quantum entanglement in an experimentally feasible dissipative optomechanical system Sawadsky et al.
@@
 (2011).
-The setup consists of a fixed perfect mirror and an effective movable mirror, realized by a Michelson–Sagnac interferometer (MSI) incorporating a membrane.
-Compared to dispersive optomechanics Vitali et al.
-(2007) with negligible dissipative coupling, the dispersive and dissipative couplings can be independently switched on and off by adjusting the membrane position or the reflectivity (transmissivity) of the beam splitter (BS) in this unconventional dissipative optomechanics Xuereb et al.
+The setup consists of a fixed perfect mirror and an effectively movable mirror, realized by a Michelson–Sagnac interferometer (MSI) with an integrated membrane.
+We begin by analyzing the steady-state mechanical displacement under both coherent and dissipative coupling.
+Notably, with coherent coupling, the mechanical displacement exhibits a nonlinear dependence on the driving power, consistent with conventional optomechanics Aspelmeyer et al.
 (2014).
-We first examine the behavior of the steady-state mechanical displacement with purely dispersive or dissipative coupling.
-It is evident that that the displacement exhibits nonlinear dependence on the driving power with purely dispersive coupling, similar to conventional optomechanical systems Aspelmeyer et al.
-(2014).
-However, it becomes linear dependence of the driving power for purely dissipative coupling.
-We then study the effect of the purely dispersive or dissipative coupling on generation of quantum entanglement.
-We surprisely find that the quantum entanglement induced by purely dissipative coupling is not only much stronger than that generated by purely dispersive coupling, but also it is more robust aginst the bath temperature.
-For purely dispersive entanglement, its survival temperature is approximately $\sim 8.5\,\text{K}$, while for purely dissipative entanglement, it can be up to $\sim 25\,\text{K}$.
-When both couplings coexist, entanglement reduces due to quantum interference.
-Our work provides a promising path to generate strong and noise-resilient quantum entanglement in quantum systems with purely dissipative coupling.
+In contrast, under dissipative coupling, the displacement shows a linear dependence on the driving power.
+Furthermore, we explore quantum entanglement in the presence of coherent and dissipative coupling.
+Interestingly, we find that entanglement induced by dissipative coupling is not only significantly stronger than that generated by coherent coupling but also remarkably more robust against bath temperature.
+For entanglement with coherent optomechanical coupling, its survival temperature is approximately $\sim 8.5\,\text{K}$, while for entanglement with dissipative optomechanical coupling, it can be up to $\sim 25\,\text{K}$.
+When both the coherent and dissipative couplings are simultaneously considered, entanglement instead reduces due to quantum interference.
+Our work provides a promising path to generate strong and noise-resilient quantum entanglement in quantum systems with dissipative coupling.
 The rest paper is organized as follows.
@@
 III Steady-state solution
-When dissipation is included, the dynamics of the proposed system can be governed by the quantum Langevin equantion $\mathcal{O}=-i[\mathcal{O},\mathcal{H}_{S}]+{Dissipation}+{Noise}$, specifically,
+When dissipation is included, the dynamics of the proposed system can be governed by the quantum Langevin equantion $\dot{\mathcal{O}}=-i[\mathcal{O},\mathcal{H}_{S}]+{Dissipation}+{Noise}$, specifically,
 $\dot{x}=$ $\omega_{m}p,$
@@
 Figure 2: The normalized steady-state displacment $x_{s}$ of the membrane vs (a) the intracavity mean photon number $n_{s}=|a_{s}|^{2}$ and (b) the power of the driving field.
-The red curve corresponds to the case of the purely dispersive coupling ($g_{\omega}/2\pi=2$ Hz).
-The blue curve is plotted for the purely dissipative coupling ($g_{\kappa}/2\pi=2$ Hz).
+The red curve corresponds to the case of the coherent coupling ($g_{\omega}/2\pi=2$ Hz).
+The blue curve is plotted for the dissipative coupling ($g_{\kappa}/2\pi=2$ Hz).
 In (a) and (b), $\Delta_{a}=3\kappa_{a}$ is assumed.
@@
 Obivously, the steady-state displacement of the membrane is determined by both the dispersive and dissipative coupling between photons and phonons.
-For a purely dispersive coupling ($g_{\kappa}=0$), $x_{s}$ is proportional to the intracavity mean photon number $n_{s}=|a_{s}|^{2}$ [see the red curve in Fig.
+For a coherent coupling ($g_{\kappa}=0$), $x_{s}$ is proportional to the intracavity mean photon number $n_{s}=|a_{s}|^{2}$ [see the red curve in Fig.
 2(a)].
-For a purely dissipative coupling ($g_{\omega}=0$), $x_{s}$ becomes
+For a dissipative coupling ($g_{\omega}=0$), $x_{s}$ becomes
 $x_{s}=[1-\sqrt{1-2(g_{\kappa}/\kappa_{a})^{2}(\Delta_{a}/\omega_{m})n_{s}}]\omega_{m}\kappa_{a},$ (10)
@@
 When $n_{s}=0$, $x_{s}=0$, i.e., the membrane is positioned at its equilibrum.
-Here $x_{s}$ and $n_{s}$ are quadratic dependence for the purely dissipative coupling, demonstrated by the blue curve in Fig.
+Here $x_{s}$ and $n_{s}$ are quadratic dependence for the dissipative coupling, demonstrated by the blue curve in Fig.
 2(a).
-One may further notice that the displacement of the membrane caused by the purely dispersive coupling is larger than that caused by the purely dissipative coupling, as clearly shown in Fig.
+One may further notice that the displacement of the membrane caused by the coherent coupling is larger than that caused by the dissipative coupling, as clearly shown in Fig.
 2(a).
@@
 (11) may have three real roots, corresponding to bistability.
-But in fact, the bistability can only be predicted for puly dispersive coupling, as shown by the red curve in Fig.
+But in fact, the bistability can only be predicted for coherent coupling, as shown by the red curve in Fig.
 2(b).
-This indicates that the purely dissipative coupling can not generate the typical stability in optomechanics, demonstrated by the blue curve in Fig.
+This indicates that dissipative coupling can not generate the typical bistability in optomechanics, demonstrated by the blue curve in Fig.
 2(b).
@@
 (III), we then obtain a set of equations for fluctuations ($\delta\mathcal{O}$).
-For the strong driving field ($|a_{s}|^{2}\gg 1$), the higher-order fluctuation terms can be safely neglected Vitali et al.
+For the strong driving field ($n_{s}\gg 1$), the higher-order fluctuation terms can be safely neglected Vitali et al.
 (2007), corresponding to linearization.
@@
 $D=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&\gamma_{m}(2n_{\rm th}+1)+\frac{G_{\kappa}^{2}}{4\kappa_{a}}&0&\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}\frac{x_{s}^{\prime}}{2}\\ 0&0&\frac{x_{s}^{\prime,2}}{2}&0\\ 0&\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}\frac{x_{s}^{\prime}}{2}&0&\frac{x_{s}^{\prime,2}}{2}\\ \end{array}\right).$ (17)
-Compared to the case of purely dispersive coupling, the dissipative coupling strength appears in the off-diagonal elements of the diffusion matrix, rendering $D$ non-diagonal.
+Compared to coherent coupling case, dissipative coupling strength appears in the off-diagonal elements of the diffusion matrix, playing an crucial role in generation of bipartite entanglement below.
 To measure optomechanical entanglement, the logarithmic negativity ($E_{N}$) is taken Adesso et al. (2004), i.e.,
@@
 V Optomechanical Entanglement
-To investigate optomechanical entanglement, we consider three specific scenarios: (i) purely dispersive coupling, (ii) purely dissipative coupling, and (iii) the coexistence of both couplings.
-V.1 Purely dispersive coupling ($g_{\kappa}=0$)
-In the purely dispersive coupling regime ($g_{\kappa}=0$), we present the entanglement measure $E_{N}$ as a function of both the dispersive coupling strength $g_{\omega}$ and the normalized effective cavity detuning $\Delta_{s}/\omega_{m}$ in Fig.
+To investigate optomechanical entanglement, we consider three specific scenarios: (i) coherent coupling, (ii) dissipative coupling, and (iii) cooperative coupling (i.e., both the coherent and dissipative couplings are considered).
+V.1 Coherent coupling ($g_{\kappa}=0$)
+In the coherent coupling case ($g_{\kappa}=0$), we present the entanglement measure $E_{N}$ as a function of both the coherent coupling strength $g_{\omega}$ and the normalized effective cavity detuning $\Delta_{s}/\omega_{m}$ in Fig.
 3(a) to explore the conditions for optimal optomechanical entanglement at a bath temperature of $T=0.4$ K.
-The results reveal that the generated entanglement remains relatively weak compared to that typically observed in conventional optomechanical systems.
-This is primarily due to the inherent limitation in the dispersive coupling strength $g_{\omega}$ (or equivalently $G_{\omega}$), which must remain small within the stable regime to avoid system instability, thereby restricting the achievable entanglement.
+The results reveal that the generated entanglement remains relatively weak compared to that typically observed in the conventional optomechanical system Vitali et al.
+(2007).
+This is because the stronger coupling strength will lead to system unstable.
+To numerically ensure system stable, the coupling strength is required to be weak, giving rise to weak entanglement.
 Nonetheless, optimal entanglement can still be achieved around the point ($\Delta_{s}/\omega_{m},g_{\omega}/2\pi)\approx(2,3.1)$, as marked by the green star, indicating that even under these constraints, entanglement can be optimized.
-Figure 3: (a) The logarithmic negativity $E_{N}$ vs the normalized detuning $\Delta/\omega_{m}$ and the purely dispersive coupling strength $g_{\omega}$ at bath temperature $T=400$ mK.
-(b) The logarithmic negativity $E_{N}$ vs the dispersive coupling strength $g_{\omega}$ with $T=400$ mK (see the red curve) and $T=4$ K (see the blue dashed curve), where $\Delta_{s}/\omega_{m}=2$ is fixed.
+Figure 3: (a) The logarithmic negativity $E_{N}$ vs the normalized detuning $\Delta/\omega_{m}$ and the coherent coupling strength $g_{\omega}$ at bath temperature $T=400$ mK.
+(b) The logarithmic negativity $E_{N}$ vs the coherent coupling strength $g_{\omega}$ with $T=400$ mK (see the red curve) and $T=4$ K (see the blue dashed curve), where $\Delta_{s}/\omega_{m}=2$ is fixed.
 (c) The logarithmic negativity $E_{N}$ vs the bath temperature $T$ with $g_{\omega}/2\pi=3.1$ Hz and $\Delta_{s}/\omega_{m}=2$.
@@
 2 except for $\mathcal{P}=50$ mW.
-To further investigate the role of dispersive coupling strength on entanglement, we plot $E_{N}$ as a function of $g_{\omega}$ in Fig.
+To further investigate the coherent coupling effect on entanglement, we plot $E_{N}$ as a function of $g_{\omega}$ in Fig.
 3(b), where $\Delta_{s}\approx 2\omega_{m}$ is fixed.
@@
 When the bath temperature increases to $T\approx 8.5$ K, $E_{N}=0$, that is, quantum entanglment vanishes and the system becomes separable.
-V.2 Purely dissipative coupling ($g_{\omega}=0$)
-Figure 4: (a) The logarithmic negativity $E_{N}$ vs the normalized detuning $\Delta/\omega_{m}$ and the purely dissipative coupling strength $g_{\kappa}$ at bath temperature $T=400$ mK.
+V.2 Dissipative coupling ($g_{\omega}=0$)
+In dissipative coupling case ($g_{\omega}=0$), we observe that the optimal entanglement can be significantly stronger than in the coherent coupling case, even surpassing the entanglement achieved in conventional optomechanical system (i.e., $E_{N}=0.35$Vitali et al.
+(2007).
+This is demonstrated in Fig.
+4(a), where $E_{N}$ is plotted as a function of both the dissipative coupling strength ($g_{\kappa}$) and the normalized effective cavity detuning $\Delta_{s}/\omega_{m}$.
+The optimal entanglement is found at approximately ($\Delta_{s}/\omega_{m},g_{\kappa}/2\pi)\approx(0.1,20)$, as indicated by the green star.
+In contrast to the coherent coupling case [see Fig.
+2(a)], where the optimal entanglement occurs near the mechanical red sideband at $\Delta_{s}/\omega_{m}=2$, the optimal entanglement in the dissipative coupling case is obtained when the cavity is nearly resonant with the driving field, specifically at $\Delta_{s}/\omega_{m}=0.1$, which is far removed from the mechanical red sideband.
+Figure 4: (a) The logarithmic negativity $E_{N}$ vs the normalized detuning $\Delta/\omega_{m}$ and the dissipative coupling strength $g_{\kappa}$ at bath temperature $T=400$ mK.
 (b) The logarithmic negativity $E_{N}$ vs the dissipative coupling strength $g_{\kappa}$ with $T=400$ mK (see the red curve) and $T=4$ K (see the blue dashed curve), where $\Delta_{s}/\omega_{m}=0.1$ is fixed.
@@
 2 except for $\mathcal{P}=50$ mW.
-In the purely dissipative case ($g_{\omega}=0$), we observe that the optimal entanglement can be significantly stronger than in the purely dispersive case, even surpassing the entanglement achieved in conventional optomechanical systems (i.e., $E_{N}=0.35$).
-This is demonstrated in Fig.
-4(a), where $E_{N}$ is plotted as a function of both the dissipative coupling strength ($g_{\kappa}$) and the normalized effective cavity detuning $\Delta_{s}/\omega_{m}$.
-The optimal entanglement is found at approximately ($\Delta_{s}/\omega_{m},g_{\kappa}/2\pi)\approx(0.1,20)$, as indicated by the green star.
-In contrast to the purely dispersive case [see Fig.
-2(a)], where the optimal entanglement occurs near the mechanical red sideband at $\Delta_{s}/\omega_{m}=2$, the optimal entanglement in the purely dissipative case is obtained when the cavity is nearly resonant with the driving field, specifically at $\Delta_{s}/\omega_{m}=0.1$, which is far removed from the mechanical red sideband.
 Furthermore, in Fig.
 4(b), we investigate the behavior of the optomechanical entanglement as a function of the dissipative coupling strength, while fixing $\Delta_{s}=0.1\omega_{m}$.
-The resulting behavior of $E_{N}$ mirrors that observed in the purely dispersive coupling case (see the red curve).
+The resulting behavior of $E_{N}$ mirrors that observed in the coherent coupling case (see the red curve).
 The abrupt drop in entanglement marks the transition of the system from a stable to an unstable region, as indicated by the pink area.
@@
 However, beyond this point, the entanglement becomes more tolerant to further temperature increases, suggesting that the dissipative coupling becomes more robust to thermal noise at higher temperatures.
-Remarkably, entanglement persists even up to 25 K, This indicates that strong, noise-resistant entanglement can be generated through purely dissipative coupling, opening new possibilities for quantum information processing and communication in noisy
+Remarkably, entanglement persists even up to 25 K, This indicates that strong, noise-resistant entanglement can be generated through dissipative optomechanical coupling, opening new possibilities for quantum information processing and communication in noisy
+V.3 Cooperative coupling ($g_{\omega}\neq 0$ and $g_{\kappa}\neq 0$)
 Figure 5: (a) The logarithmic negativity $E_{N}$ vs the normalized detuning $\Delta/\omega_{m}$ and the ratio of $g_{\kappa}/g_{\omega}$ at bath temperature $T=400$ mK.
@@
 2 except for $\mathcal{P}=50$ mW, $g_{\omega}/2\pi=3$ Hz.
-V.3 Combined coupling ($g_{\omega}\neq 0$ and $g_{\kappa}\neq 0$)
-To examine the interplay between dispersive and dissipative coupling effects on optomechanical entanglement, we set the dispersive coupling strength to $g_{\omega}/2\pi=3$ Hz, corresponding to the optimal entanglement condition at $\Delta_{s}/\omega_{m}=2$ in the purely dispersive coupling regime [refer to Fig.
+To examine the effect of the cooperative coupling (i.e., coherent and dissipative couplings) on optomechanical entanglement, we set the coherent coupling strength to be $g_{\omega}/2\pi=3$ Hz, corresponding to the optimal entanglement condition at $\Delta_{s}/\omega_{m}=2$ in the dispersive coupling case [refer to Fig.
 3(a)].
@@
 5(a).
-The reduction in entanglement possibly arises from the interference between the dispersive and dissipative couplings.
+The reduction in entanglement possibly arises from the interference between the coherent and dissipative couplings.
 When the bath temperature increases to $T=4$ K, it is observed that the optomechancial entanglement is further suppressed, as depicted by the blue dashed curve in Fig.
@@
 VI Conclusion
-In this paper, we investigate quantum entanglement in a dissipative optomechanical system, acted by a fixed mirror and a Michelson–Sagnac interferometer with a membrane.
-This system allows the dissipative and dispersive couplings to be independently tuned on and off by adjusting the membrane position or the reflectivity (transmissivity) of the BS, enabling us to explore quantum entanglement in three distinct scenarios: purely dispersive case, purely dissipative case, and combined coupling case.
-Our analysis reveals that, in the purely dispersive case, the steady-state mechanical displacement exhibits a nonlinear dependence on the driving power, whereas in the purely dissipative case, the displacement shows a linear dependence.
-Furthermore, entanglement generated through purely dissipative coupling is not only significantly stronger but also more noise-tolerant compared to that achieved via purely dispersive coupling.
+In a summary, we investigate quantum entanglement in a dissipative optomechanical system, realized by a fixed mirror and a Michelson–Sagnac interferometer with a membrane.
+This system allows the dissipative and coherent couplings to be independently tuned on and off by adjusting the membrane position or the reflectivity (transmissivity) of the beam splitter, enabling us to explore quantum entanglement in three distinct scenarios: coherent coupling case, dissipative case, and cooperative case.
+Our results reveals that, in the coherent coupling case, the steady-state mechanical displacement exhibits a nonlinear dependence on the driving power, whereas in the dissipative coupling case, the displacement shows a linear dependence.
+Furthermore, entanglement generated through dissipative coupling is not only significantly stronger but also more noise-tolerant compared to that achieved via coherent coupling.
 When both coupling mechanisms coexist, quantum entanglement is reduced, likely due to the interference between the two coupling interactions.
-These findings suggest the potential of purely dissipative coupling to engineer strong and noise-tolerant quantum entanglement, offering a promising pathway for advancing quantum technologies.
+These findings suggest that dissipative coupling may be a promising mechanism to engineer strong and noise-tolerant quantum entanglement.
 This work is supported by the National Natural Science Foundation of China (Grant No.

### v4

@@
 Abstract
-Macroscopic entanglement, as critical quantum resources in quantum information science, has been extensively studied in coherent optomechanics over the past decades.
-However, entanglement in dissipative optomechanics, where the cavity linewidth is dependent on the position of the mechanical resonator, remains unexplored.
-In this work, we study quantum entanglement in a dissipative optomechanics, realized by a Michelson–Sagnac interferometer and a movable membrane.
-Such dissipative optomechanics allows to switch coherent and dissipative optomechanical coupling at will.
-With reliable parameters, we demonstrate that the steady-state mechanical displacement exhibits a nonlinear (linear) dependence on the driving power with coherent (dissipative) coupling.
-Further, we show that quantum entanglement generated by dissipative coupling is significantly stronger and more noise-tolerant than that generated by coherent coupling.
-When both the coherent and dissipative couplings are simultaneously considered, entanglement is instead weakened due to quantum interference.
-Our result indicates that dissipative optomechanical coupling may be a promising path to engineer strong and noise-tolerant quantum entanglement.
+Macroscopic entanglement, as a critical quantum resource in quantum information science, has been extensively studied in coherent optomechanics over the past decades.
+However, entanglement in dissipative optomechanics, where the cavity linewidth depends on the position of the mechanical resonator, remains largely unexplored.
+In this work, we investigate quantum entanglement in a dissipative optomechanical system realized by a Michelson-Sagnac interferometer with a movable membrane.
+This configuration enables the switching between coherent and dissipative optomechanical couplings at will.
+With experimentally feasible parameters, we demonstrate that the steady-state mechanical displacement exhibits a nonlinear (linear) dependence on the driving power under coherent (dissipative) coupling.
+Furthermore, we show that the quantum entanglement generated via dissipative coupling is significantly stronger and more robust to noise than that generated via coherent coupling.
+When both coherent and dissipative couplings are simultaneously present, the entanglement is weakened due to quantum interference.
+Our results indicate that dissipative optomechanical coupling can be a promising route for engineering strong and noise-resilient quantum entanglement.
 I Introduction
-Coherent optomechanics, where the cavity frequency is dependent on the position of the mechanical resonator, has garnered siginificant interest theoretically and experimentally over the past two decades Aspelmeyer et al.
+Coherent optomechanics, where the cavity frequency depends on the position of the mechanical resonator, has garnered significant interest both theoretically and experimentally over the past two decades Aspelmeyer et al.
 (2014).
@@
 (1985) and tristability Xiong et al.
-(2016), optomechancally induced transparency Weis et al.
+(2016), optomechanically induced transparency Weis et al.
 (2010); Safavi-Naeini et al.
@@
 (2013); Luo et al.
-(2022), phase transition Wang et al.
+(2022), phase transitions Wang et al.
 (2024); Jäger et al.
@@
 (2023a); Xu et al.
-(2016), spin-optomechanical interface Xiong et al.
+(2016), spin-optomechanical interfaces Xiong et al.
 (2021a); Chen et al.
@@
 (2023); Jäger et al.
-(2019), higher-order exceptional point Xiong et al.
+(2019), higher-order exceptional points Xiong et al.
 (2021b, 2022), and macroscopic quantum entanglement Vitali et al.
@@
 (2009), giving rise to a dissipative coupling between the mechanical and cavity modes.
-Later, the setup was extended to a Michelson-Sagnac interferometers (MSI) with a movable membrane Xuereb et al.
-(2011), where coherent or dissipative optomechanical coupling is allowed to switch on and off at will by tuning the position of the membrane or the reflectivity of the beam splitter.
-Experimentally, dissipative optomechanics have been demonstrated in MSI Sawadsky et al.
+Later, the setup was extended to Michelson-Sagnac interferometers (MSIs) with a movable membrane Xuereb et al.
+(2011); Tarabrin et al.
+(2013), where coherent or dissipative optomechanical coupling can be switched on and off at will by tuning the position of the membrane or the reflectivity amplitude of the beam splitter (BS).
+Experimentally, dissipative optomechanics has been demonstrated in MSIs Sawadsky et al.
 (2015); Friedrich et al.
-(2011), microdisc resonator coupled to a nanomechanical waveguide Li et al.
-(2009), photonic crystal split-beam nanocavity Wu et al.
-(2014), graphene drums coupled to high-Q microsphere Cole et al.
-(2015), and levitated sphere trapped by dual-beam optical tweezers Kuang et al.
+(2011), a microdisk resonator coupled to a nanomechanical waveguide Li et al.
+(2009), a photonic crystal split-beam nanocavity Wu et al.
+(2014), graphene drums coupled to a high-Q microsphere Cole et al.
+(2015), and a levitated sphere trapped by dual-beam optical tweezers Kuang et al.
 (2023).
@@
 (2023); Zhang et al.
-(2022) have been investigated.
-However, macroscopic quantum entanglement has not been revealed yet with dissipative optomechanical coupling, although it has been intensively studied in coherent optomechanics.
-In this work, we propose to investigate quantum entanglement in an experimentally feasible dissipative optomechanical system Sawadsky et al.
+(2022), have been investigated.
+However, macroscopic quantum entanglement has not yet been revealed in dissipative optomechanical coupling, although it has been intensively studied in coherent optomechanics.
+In this work, we investigate quantum entanglement in an experimentally realizable dissipative optomechanical system Sawadsky et al.
 (2015); Friedrich et al.
 (2011).
-The setup consists of a fixed perfect mirror and an effectively movable mirror, realized by a Michelson–Sagnac interferometer (MSI) with an integrated membrane.
+The setup consists of a fixed perfect mirror and an effectively movable mirror implemented using a Michelson-Sagnac interferometer (MSI) with an integrated membrane.
 We begin by analyzing the steady-state mechanical displacement under both coherent and dissipative coupling.
-Notably, with coherent coupling, the mechanical displacement exhibits a nonlinear dependence on the driving power, consistent with conventional optomechanics Aspelmeyer et al.
+Notably, with coherent coupling, the mechanical displacement exhibits a nonlinear relationship with the driving power, consistent with conventional optomechanics Aspelmeyer et al.
 (2014).
 In contrast, under dissipative coupling, the displacement shows a linear dependence on the driving power.
-Furthermore, we explore quantum entanglement in the presence of coherent and dissipative coupling.
-Interestingly, we find that entanglement induced by dissipative coupling is not only significantly stronger than that generated by coherent coupling but also remarkably more robust against bath temperature.
-For entanglement with coherent optomechanical coupling, its survival temperature is approximately $\sim 8.5\,\text{K}$, while for entanglement with dissipative optomechanical coupling, it can be up to $\sim 25\,\text{K}$.
-When both the coherent and dissipative couplings are simultaneously considered, entanglement instead reduces due to quantum interference.
-Our work provides a promising path to generate strong and noise-resilient quantum entanglement in quantum systems with dissipative coupling.
+Furthermore, we explore quantum entanglement in the presence of both coherent and dissipative coupling.
+Interestingly, we find that entanglement induced by dissipative coupling is not only significantly stronger than that generated by coherent coupling but also considerably more robust against bath temperature.
+For entanglement with coherent optomechanical coupling, its survival temperature is approximately $\sim 8.5$ K, while for entanglement with dissipative optomechanical coupling, it can reach up to $\sim 25$ K.
+When both coherent and dissipative couplings are simultaneously considered, entanglement is reduced due to quantum interference.
+Our work provides a promising path to generate strong and temperature-resilient quantum entanglement in quantum systems with dissipative coupling.
 The rest paper is organized as follows.
@@
 II The model and Hamiltonian
-Figure 1: (a) Schematic diagram of a typical dissipative optomechanical system, realized by a perfect fixed mirror $M_{1}$ and an effective movable mirror $\mathcal{M}$.
-(b) The structure of the effective movable mirror $\mathcal{M}$.
-The effective mirror consists of a Michelson-Sagnac interferometer (made up of a beam-splitter and two mirrors $M_{2}$ and $M_{3}$) and a movable membrane (mechanical resonator).
-$L$ is the effective length of the cavity ($2L$ is the length of the Sagnac mode ${\rm M_{1}-BS-M_{2}-M_{3}-BS}$).
-$a_{\rm in}$ is the input port and $a_{\rm out}$ is the output port.
+Figure 1: (a) Schematic diagram of a typical dissipative optomechanical system, consisting of a perfectly fixed mirror $M_{1}$ and an effective movable mirror $\mathcal{M}$.
+(b) Structure of the effective movable mirror $\mathcal{M}$.
+It comprises a Michelson-Sagnac interferometer, which is formed by a beam splitter and two mirrors $M_{2}$ and $M_{3}$, along with a movable membrane (mechanical resonator).
+The parameter $L$ denotes the effective cavity length, where $2L$ is the round-trip length of the Sagnac mode ${\rm M}_{1}$-${\rm BS}$-${\rm M}_{2}$-${\rm M}_{3}$-${\rm BS}$.
+The optical field enters the system through the input port $a_{\rm in}$ and exits through the output port $a_{\rm out}$.
 As shown in Fig.
-1(a), we consider an experimentally feasible dissipative optomechanics Sawadsky et al.
+1(a), we consider an experimentally feasible dissipative optomechanical system Sawadsky et al.
 (2015); Friedrich et al.
 (2011), consisting of a fixed mirror $M_{1}$ and an effective movable mirror $\mathcal{M}$.
-The effective mirror can be realized by a Michelson-Sagnac interferometer (MSI) and a movable membrane between mirrors $M_{2}$ and $M_{3}$ [see Fig.
+The effective mirror can be realized by a Michelson-Sagnac interferometer (MSI) incorporating a movable membrane placed between mirrors $M_{2}$ and $M_{3}$ [see Fig.
 1(b)].
-$L$ is the effective length of the cavity ($2L$ is the length of the Sagnac mode ${\rm M}_{1}-{\rm BS}-{\rm M}_{2}-{\rm M}_{3}-{\rm BS}$).
-The membrane is treated quantum mechanically, with its displacement and momentum operators denoted by $x$ and $p$, respectively, satisfying the commutation relation $[x,p]=i\hbar$.
-Upon the introduction of an external driving field (with frequency $\omega_{d}$ and power $P$) at the beam splitter, the system can be effectively modeled as a typical dissipative optomechanical system.
-Thus, both the cavity frequency $\omega_{a}(x)$ and the decay rate $\kappa_{a}(x)$ depend on the displacement $x$ of the membrane.
-For small membrane displacements $x$, we can Taylor expand $\omega_{a}(x)$ and $\kappa_{a}(x)$ to the first order, i.e., $\omega(x)\approx\omega_{a}-g_{\omega}x$ and $\kappa(x)\approx\kappa_{a}-g_{\kappa}x$, where $\omega_{a}$ ($\kappa_{a}$) is the frequency (decay rate) of the cavity with annihilation (creation) operator $a$ ($a^{\dagger}$) when the membrane is at its equilibrium.
-The Hamiltonian of the considered system is given by Elste et al.
+Here, $L$ denotes the effective length of the cavity (with $2L$ representing the round-trip length of the Sagnac mode ${\rm M}_{1}$-${\rm BS}$-${\rm M}_{2}$-${\rm M}_{3}$-${\rm BS}$).
+The membrane is treated quantum mechanically, with its displacement and momentum operators denoted by $x$ and $p$, respectively, satisfying the canonical commutation relation $[x,p]=i\hbar$.
+Upon the introduction of an external driving field (with frequency $\omega_{d}$ and power $P$) at BS, the system can be effectively described as a typical dissipative optomechanical system, where both the cavity resonance frequency $\omega_{a}(x)$ and the decay rate $\kappa_{a}(x)$ depend on the membrane displacement $x$.
+In the good-cavity limit Yamamoto et al.
+(2010), the system Hamiltonian is given by Elste et al.
 (2009); Xuereb et al.
-(2011)
+(2011); Tarabrin et al.
+(2013) (setting $\hbar=1$)
 $H_{S}=$ $(\omega_{a}-g_{\omega}x)a^{\dagger}a+\frac{1}{2}\omega_{m}(x^{2}+p^{2})$
 […]
-where $\omega_{m}$ is the frequency of the membrane, $\omega$ is the frequency of the bosonic mode (i.e., bath) with annihilation (creation) operator $a_{\omega}$ ($a_{\omega}^{\dagger}$), $x^{\prime}=\sqrt{2\kappa_{a}}-{g_{\kappa}x}/{\sqrt{2\kappa_{a}}}$ is the position-dependent coupling strength between the optical cavity and the bosonic mode.
-In fact, the bosonic mode here is to introduce a zero-value expectation input field $a_{\rm in}=\int\frac{d\omega}{\sqrt{2\pi}}e^{-i\omega(t-t_{0})}a_{\omega,0}$, with $a_{\omega,0}$ the operator at an initial time, to the cavity mode $a$, which can be easily obtained following the standard input-output theory with the Markov approximation.
+where $\omega_{a}$ ($\kappa_{a}$) denotes the cavity resonance frequency (decay rate) with annihilation (creation) operator $a$ ($a^{\dagger}$) when the membrane is at equilibrium, $\omega_{m}$ is the mechanical resonance frequency of the membrane, and $\omega$ is the frequency of the bosonic bath mode with annihilation (creation) operator $a_{\omega}$ ($a_{\omega}^{\dagger}$).
+The position-dependent coupling strength between the optical cavity and the bath is given by ${\Gamma}=\sqrt{2\kappa_{a}}-{g_{\kappa}x}/{\sqrt{2\kappa_{a}}}$.
+Here, $g_{\omega(\kappa)}$ represents the single-photon coherent (dissipative) optomechanical coupling strength, explicitly expressed as Elste et al.
+(2009); Xuereb et al.
+(2011); Tarabrin et al.
+(2013)
+$g_{\omega}=$ $-2(\omega_{a}x_{\rm zpf}/L)[(|R|^{2}-|T|^{2})+\tau\cos(\arg\mathcal{T})],$
+$g_{\kappa}=$ $-i\sqrt{2}(\omega_{a}x_{\rm zpf}/L)|\tau|[2RT+\rho\cos(\arg\mathcal{T})],$ (2)
+with
+$\rho=$ $-[(R^{2}e^{2ix}+T^{2}e^{-2ix})\mathcal{R}+2RT\mathcal{T}]e^{-i\rm{arg}\mathcal{T}},$
+$\tau=$ $[(RT^{*}e^{2ix}-{\rm c.c.})\mathcal{R}-(|R|^{2}-|T|^{2})\mathcal{T}]e^{-i\rm{arg}\mathcal{T}}$ (3)
+being the effective complex reflectivity and transmissivity of the mirror $\mathcal{M}$, which depend on the membrane’s position $x$.
+Here, $x_{\rm zpf}$ denotes the zero-point fluctuation of the membrane, and $R~{}(\mathcal{R})$ and $T~{}(\mathcal{T})$ represent the complex reflectivity and transmissivity of BS (membrane), respectively.
+Therefore, the values of $g_{\omega}$ and $g_{\kappa}$ can be independently tuned, allowing one to switch them on or off as desired.
+In particular, a purely dissipative coupling (i.e., $g_{\omega}=0$) can be achieved by ensuring $|R|\neq|T|$.
+Note that the bosonic mode $a_{\omega}$ in Eq.
+(II) is to introduce a zero-value expectation input field $a_{\rm in}=\int\frac{d\omega}{\sqrt{2\pi}}e^{-i\omega(t-t_{0})}a_{\omega,0}$, with $a_{\omega,0}$ the operator at an initial time, to the cavity mode $a$, which can be easily obtained following the standard input-output theory with the Markov approximation.
 At the optical domain, the input field satisfies the relations Walls and Milburn (1994)
@@
 $\mathcal{H}_{S}=$ $(\Delta_{a}-g_{\omega}x)a^{\dagger}a+\frac{1}{2}\omega_{m}(x^{2}+p^{2})$
-$+ix^{\prime}(A_{\rm in}a^{\dagger}-A_{\rm in}^{\dagger}a),$ (3)
+$+i{\Gamma}(A_{\rm in}a^{\dagger}-A_{\rm in}^{\dagger}a),$ (5)
 where $\Delta_{a}=\omega_{a}-\omega_{d}$ is the frequency detuning of the cavity mode from the driving field, and
@@
 $\dot{p}=$ $g_{\omega}a^{\dagger}a-\omega_{m}x+\frac{ig_{\kappa}}{\sqrt{2\kappa_{a}}}(A_{\rm in}^{\dagger}a-A_{\rm in}a^{\dagger})-\gamma_{m}p+\xi,$
-$\dot{a}=$ $-(\kappa_{a}+i\Delta_{a})a+ig_{\omega}ax+g_{\kappa}ax+x^{\prime}A_{\rm in},$ (5)
+$\dot{a}=$ $-(\kappa_{a}+i\Delta_{a})a+ig_{\omega}ax+g_{\kappa}ax+{\Gamma}A_{\rm in},$ (7)
 where $\gamma_{m}$ is the damping rate of the membrane and $\xi$ is the Brownian noise operator arising from the coupling of the membrane oscillator to the hot environment.
-The correlation function of $\xi$ is given by Giovannetti and Vitali (2001) $\langle\xi(t)\xi(t^{\prime})\rangle=\frac{\gamma_{m}}{\omega_{m}}\int\frac{d\omega}{2\pi}e^{-i\omega(t-t^{\prime})}\omega[{\rm coth}(\frac{\hbar\omega}{2k_{b}T})+1]$.
-In the limit of the high quality factor, $\omega_{m}\gg\gamma_{m}$, the Brownian noise $\xi$ becomes am approximate Markovian noise Benguria and Kac (1981), i.e.,
+Since the Brownian noise is inherently random and unbiased.
+In the statistical sense, its average value is zero, i.e., $\langle\xi\rangle=0$.
+However, investigating the fluctuation behavior of the system, the contribution from the correlation function of $\xi$ should be taken into account, given by Giovannetti and Vitali (2001) $\langle\xi(t)\xi(t^{\prime})\rangle=\frac{\gamma_{m}}{\omega_{m}}\int\frac{d\omega}{2\pi}e^{-i\omega(t-t^{\prime})}\omega[{\rm coth}(\frac{\hbar\omega}{2k_{b}T})+1]$.
+In the limit of the high quality factor, $\omega_{m}\gg\gamma_{m}$, the Brownian noise $\xi$ approximately becomes a Markovian noise Benguria and Kac (1981), i.e.,
 $\langle\xi(t)\xi(t^{\prime})+\xi(t^{\prime})\xi(t)\rangle/2\backsimeq\gamma_{m}(2n_{\rm th}+1)\delta(t-t^{\prime})$ (8)
@@
 In (a) and (b), $\Delta_{a}=3\kappa_{a}$ is assumed.
-When the evolution time is long enough, the system reaches its steady state, denoted by $a_{s}$, $x_{s}$, and $p_{s}$. From Eq. (III), the steady-state equation can be directly obtained as
+In the long-time limit ($t\rightarrow\infty$), the system reaches a steady state.
+Each operator $\mathcal{O}$ is then expanded as the sum of its classical steady-state value and a small quantum fluctuation, i.e., $\mathcal{O}=\mathcal{O}_{s}+\delta\mathcal{O}$.
+Substituting this decomposition into Eq.
+(III), we derive a set of equations that govern the steady-state values $\mathcal{O}_{s}$ as follows:
 $\dot{x}_{s}=$ $\omega_{m}p_{s},$
 $\dot{p}_{s}=$ $g_{\omega}|a_{s}|^{2}-\omega_{m}x_{s}+\frac{ig_{\kappa}}{\sqrt{2\kappa_{a}}}(\mathcal{E}^{*}a_{s}-\mathcal{E}a_{s}^{*})-\gamma_{m}p_{s},$
-$\dot{a}_{s}=$ $-(\kappa_{s}+i{\Delta}_{s})a_{s}+x_{s}^{\prime}\mathcal{E},$ (7)
-where ${\Delta}_{s}=\Delta_{a}-{g_{\omega}}x_{s}$, $\kappa_{s}=\kappa_{a}-g_{\kappa}x_{s}$, and $x_{s}^{\prime}=\sqrt{2\kappa_{a}}-{g_{\kappa}x_{s}}/{\sqrt{2\kappa_{a}}}$.
-At the steady state, we have $\dot{\mathcal{O}}_{s}=0$, equivalently
-$\omega_{m}p_{s}=$ $0,$
-$g_{\omega}|a_{s}|^{2}-\omega_{m}x_{s}+\frac{ig_{\kappa}}{\sqrt{2\kappa_{a}}}(\mathcal{E}^{*}a_{s}-\mathcal{E}a_{s}^{*})-\gamma_{m}p_{s}=$ $0,$ (8)
-$-(\kappa_{s}+i{\Delta}_{s})a_{s}+x_{s}^{\prime}\mathcal{E}=$ $0$
-By solving these equations, we have
-$p_{s}=$ $0,~{}~{}a_{s}=\frac{x_{s}^{\prime}\mathcal{E}}{\kappa_{s}+i{\Delta}_{s}},$
+$\dot{a}_{s}=$ $-(\kappa_{s}+i{\Delta}_{s})a_{s}+{\Gamma_{s}}\mathcal{E},$ (9)
+where ${\Delta}_{s}=\Delta_{a}-{g_{\omega}}x_{s}$, $\kappa_{s}=\kappa_{a}-g_{\kappa}x_{s}$, and ${\Gamma_{s}}=\sqrt{2\kappa_{a}}-{g_{\kappa}x_{s}}/{\sqrt{2\kappa_{a}}}$.
+At the steady state, we have $\dot{\mathcal{O}}_{s}=0$, yielding to
+$p_{s}=$ $0,~{}~{}a_{s}=\frac{{\Gamma_{s}}\mathcal{E}}{\kappa_{s}+i{\Delta}_{s}},$
 $x_{s}=$ $\frac{g_{\omega}n_{s}}{\omega_{m}}+\frac{ig_{\kappa}}{\sqrt{2\kappa_{a}}}\cdot\frac{a_{s}\mathcal{E}^{*}-a_{s}^{*}\mathcal{E}}{\omega_{m}}.$ (10)
-Obivously, the steady-state displacement of the membrane is determined by both the dispersive and dissipative coupling between photons and phonons.
-For a coherent coupling ($g_{\kappa}=0$), $x_{s}$ is proportional to the intracavity mean photon number $n_{s}=|a_{s}|^{2}$ [see the red curve in Fig.
+Obviously, the steady-state displacement of the membrane is determined by both the dispersive and dissipative couplings between photons and phonons.
+For coherent coupling ($g_{\kappa}=0$), $x_{s}$ is proportional to the intracavity mean photon number $n_{s}=|a_{s}|^{2}$ [see the red curve in Fig.
 2(a)].
-For a dissipative coupling ($g_{\omega}=0$), $x_{s}$ becomes
+For dissipative coupling ($g_{\omega}=0$), $x_{s}$ becomes
 $x_{s}=[1-\sqrt{1-2(g_{\kappa}/\kappa_{a})^{2}(\Delta_{a}/\omega_{m})n_{s}}]\omega_{m}\kappa_{a},$ (11)
@@
 When $n_{s}=0$, $x_{s}=0$, i.e., the membrane is positioned at its equilibrum.
-Here $x_{s}$ and $n_{s}$ are quadratic dependence for the dissipative coupling, demonstrated by the blue curve in Fig.
+Here $x_{s}$ exhibits a quadratic dependence on $n_{s}$ in the dissipative coupling regime, as demonstrated by the blue curve in Fig.
 2(a).
-One may further notice that the displacement of the membrane caused by the coherent coupling is larger than that caused by the dissipative coupling, as clearly shown in Fig.
+One can further observe that the displacement of the membrane induced by coherent coupling is larger than that induced by dissipative coupling, as clearly shown in Fig.
 2(a).
 To plot Fig.
-2(a), we adopt typical experimental parameters Sawadsky et al.
+2(a), we employ typical experimental parameters Sawadsky et al.
 (2015): the wavelength of the driving field $\lambda=1064$ nm, $\omega_{d}/2\pi=c/\lambda=281.96$ THz, the quality factor of the membrane $\mathcal{Q}_{m}=5.8\times 10^{5}$, $\omega_{m}/2\pi=136$ kHz, $\gamma_{m}/2\pi=0.23$ Hz, $L=8.7$ cm, $m=80$ ng, and $\kappa_{a}/2\pi=1.5$ MHz.
-These experimental parameters ensure the stability of the proposed system according to Routh-Hurwitz criterion.
+These parameters guarantee the stability of the proposed system according to the Routh-Hurwitz criterion.
 From Eq. (III), one can see that $x_{s}$ actually satisfies the following cubic equation related to the amplitude of the driving field $\mathcal{E}$,
@@
 $d=$ $-4\kappa_{a}|\mathcal{E}|^{2}(g_{\omega}\kappa_{a}+g_{\kappa}\Delta_{a}).$
-Explicitly, Eq.
+Specifically, Eq.
 (12) may have three real roots, corresponding to bistability.
-But in fact, the bistability can only be predicted for coherent coupling, as shown by the red curve in Fig.
+However, bistability can only be predicted for coherent coupling, as shown by the red curve in Fig.
 2(b).
-This indicates that dissipative coupling can not generate the typical bistability in optomechanics, demonstrated by the blue curve in Fig.
+This suggests that dissipative coupling cannot generate the typical bistability in optomechanics, as demonstrated by the blue curve in Fig.
 2(b).
 IV Covariance matrix
-With steady-state values in hand, we further expand each operator $\mathcal{O}$ around its steady state, i.e., $\mathcal{O}=\delta\mathcal{O}+\mathcal{O}_{s}$.
-Substituting this expression back into Eq.
-(III), we then obtain a set of equations for fluctuations ($\delta\mathcal{O}$).
-For the strong driving field ($n_{s}\gg 1$), the higher-order fluctuation terms can be safely neglected Vitali et al.
-(2007), corresponding to linearization.
-As a result, the dynamics for fluctuations can be written as
+Apart from the steady-state values, a set of equations related to the fluctuations $\delta\mathcal{O}$ can be obtained when the expression of $\mathcal{O}=\mathcal{O}_{s}+\delta\mathcal{O}$ is substituted back into Eq.
+(III).
+Under the strong driving field ($n_{s}\gg 1$), the higher-order fluctuation terms can be safely neglected Vitali et al.
+(2007), which corresponds to the linearization of the system dynamics.
+As a result, the dynamics for the fluctuations can be written as
 $\dot{\delta x}=$ $\omega_{m}\delta p,$
@@
 $+\frac{ig_{\kappa}(\mathcal{E}^{*}\delta a-\mathcal{E}\delta a^{\dagger})}{\sqrt{2\kappa_{a}}}+\frac{i(G_{\kappa}\delta a_{\rm in}^{\dagger}-G_{\kappa}^{*}\delta a_{\rm in})}{2\sqrt{\kappa_{a}}},$ (14)
-$\dot{\delta a}=$ $-(\kappa_{s}+i\Delta_{s})\delta a+(\frac{G_{\kappa}+iG_{\omega}}{\sqrt{2}}-\frac{g_{\kappa}\mathcal{E}}{\sqrt{2\kappa_{a}}})\delta x+x_{s}^{\prime}a_{\rm in},$
+$\dot{\delta a}=$ $-(\kappa_{s}+i\Delta_{s})\delta a+(\frac{G_{\kappa}+iG_{\omega}}{\sqrt{2}}-\frac{g_{\kappa}\mathcal{E}}{\sqrt{2\kappa_{a}}})\delta x+{\Gamma_{s}}a_{\rm in},$
 where $G_{\omega(\kappa)}=\sqrt{2}g_{\omega(\kappa)}a_{s}$ is the linearized dispersive (dissipative) coupling strength, enhanced by the amplitude of the driving field. By further defining
@@
 $\dot{u}(t)=Au(t)+n(t),$ (16)
-where $u^{T}(t)=(\delta x,\delta p,\delta x_{a},\delta p_{a})$ is the vector of fluctuation operators, $n(t)$ is the vector of noise operator, given by $n^{T}(t)=(0,\xi+\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}p_{\rm in},x_{s}^{\prime}x_{\rm in},x_{s}^{\prime}p_{\rm in})$, and the drift matrix $A$ is
-$A=\left(\begin{array}[]{cccc}0&\omega_{m}&0&0\\ -\omega_{m}&-\gamma_{m}&G_{\omega}+\frac{G_{\kappa}\Delta_{s}}{\sqrt{2\kappa_{a}}x_{s}^{\prime}}&-\frac{G_{\kappa}\kappa_{s}}{\sqrt{2\kappa_{a}}x_{s}^{\prime}}\\ G_{\kappa}-\frac{2G_{\kappa}\kappa_{s}}{\sqrt{2\kappa_{a}}x_{s}^{\prime}}&0&-\kappa_{s}&\Delta_{s}\\ G_{\omega}-\frac{2G_{\kappa}\Delta_{s}}{\sqrt{2\kappa_{a}}x_{s}^{\prime}}&0&-\Delta_{s}&-\kappa_{s}\\ \end{array}\right).$ (16)
+where $u^{T}(t)=(\delta x,\delta p,\delta x_{a},\delta p_{a})$ is the vector of fluctuation operators, $n(t)$ is the vector of noise operator, given by $n^{T}(t)=(0,\xi+\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}p_{\rm in},{\Gamma_{s}}x_{\rm in},{\Gamma_{s}}p_{\rm in})$, and the drift matrix $A$ is
+$A=\left(\begin{array}[]{cccc}0&\omega_{m}&0&0\\ -\omega_{m}&-\gamma_{m}&G_{\omega}+\frac{G_{\kappa}\Delta_{s}}{\sqrt{2\kappa_{a}}{\Gamma_{s}}}&-\frac{G_{\kappa}\kappa_{s}}{\sqrt{2\kappa_{a}}{\Gamma_{s}}}\\ G_{\kappa}-\frac{2G_{\kappa}\kappa_{s}}{\sqrt{2\kappa_{a}}{\Gamma_{s}}}&0&-\kappa_{s}&\Delta_{s}\\ G_{\omega}-\frac{2G_{\kappa}\Delta_{s}}{\sqrt{2\kappa_{a}}{\Gamma_{s}}}&0&-\Delta_{s}&-\kappa_{s}\\ \end{array}\right).$ (17)
 To obtain the drift matrix $A$, we have taken $a_{s}$ as real number, which can be realized by setting $\tan\theta=\Delta_{a}/\kappa_{a}$ in Eq. (III).
 Since the phonon noise $\xi$ and the photon noise $a_{\rm in}$ are zero-mean quantum Gaussian noises and the dynamics is linearized [see Eq.
-(IV)], the quantum steady state for fluctuations is a zero-mean bipartite Gaussian state, fully described by a $4\times 4$ correlation matrix ${V}$, where $V_{ij}=\langle u_{i}(t)u_{j}(t^{\prime})+u_{j}(t^{\prime})u_{i}(t)\rangle/2$ with $i,j=1,2,3,4$.
+(IV)], the quantum steady state for fluctuations is a zero-mean bipartite Gaussian state, fully described by a $4\times 4$ correlation matrix ${V}$, where $V_{ij}=\langle u_{i}(\infty)u_{j}(\infty)+u_{j}(\infty)u_{i}(\infty)\rangle/2$ with $i,j=1,2,3,4$.
 The $V$ is obtained by solving the Lyapunov equation $AV+VA^{T}=-D$, where $D$ is the diffusion matrix, defined by $\langle n_{i}(t)n_{j}(t^{\prime})+n_{j}(t^{\prime})n_{i}(t)\rangle/2=D_{ij}\delta(t-t^{\prime})$.
 More specifically, $D$ can be expressed as
-$D=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&\gamma_{m}(2n_{\rm th}+1)+\frac{G_{\kappa}^{2}}{4\kappa_{a}}&0&\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}\frac{x_{s}^{\prime}}{2}\\ 0&0&\frac{x_{s}^{\prime,2}}{2}&0\\ 0&\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}\frac{x_{s}^{\prime}}{2}&0&\frac{x_{s}^{\prime,2}}{2}\\ \end{array}\right).$ (17)
-Compared to coherent coupling case, dissipative coupling strength appears in the off-diagonal elements of the diffusion matrix, playing an crucial role in generation of bipartite entanglement below.
+$D=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&\gamma_{m}(2n_{\rm th}+1)+\frac{G_{\kappa}^{2}}{4\kappa_{a}}&0&\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}\frac{{\Gamma_{s}}}{2}\\ 0&0&\frac{{\Gamma_{s}^{2}}}{2}&0\\ 0&\frac{G_{\kappa}}{\sqrt{2\kappa_{a}}}\frac{{\Gamma_{s}}}{2}&0&\frac{{\Gamma_{s}^{2}}}{2}\\ \end{array}\right).$ (18)
+Compared to the coherent coupling case, dissipative coupling strength appears in the off-diagonal elements of the diffusion matrix, playing a crucial role in the generation of bipartite entanglement below.
 To measure optomechanical entanglement, the logarithmic negativity ($E_{N}$) is taken Adesso et al. (2004), i.e.,
@@
 (2007).
-This is because the stronger coupling strength will lead to system unstable.
-To numerically ensure system stable, the coupling strength is required to be weak, giving rise to weak entanglement.
+This is because a weaker coupling strength is necessary to maintain system stability, which in turn leads to weaker entanglement, while a stronger coupling strength can induce instability.
 Nonetheless, optimal entanglement can still be achieved around the point ($\Delta_{s}/\omega_{m},g_{\omega}/2\pi)\approx(2,3.1)$, as marked by the green star, indicating that even under these constraints, entanglement can be optimized.
@@
 Once up to its maximum, $E_{N}$ will have a sharp reduction by further increasing $g_{\omega}$.
-Such abrupt behavior corresponds to the transition from the stability to instability, as shown by the pink-shaded region.
+Such abrupt behavior corresponds to the transition from stability to instability, as shown by the pink-shaded region.
 Additionally, when the bath temperature increases to $T=4$ K [see the blue dashed curve], the coupling for generation of quantum entanglement is required to be larger and $E_{N}$ has a significant reduction.
@@
 As the temperature increases, $E_{N}$ gradually declines.
-When the bath temperature increases to $T\approx 8.5$ K, $E_{N}=0$, that is, quantum entanglment vanishes and the system becomes separable.
+When the bath temperature increases to $T\approx 8.5$ K, $E_{N}=0$, which indicates that quantum entanglement vanishes and the system becomes separable.
 V.2 Dissipative coupling ($g_{\omega}=0$)
-In dissipative coupling case ($g_{\omega}=0$), we observe that the optimal entanglement can be significantly stronger than in the coherent coupling case, even surpassing the entanglement achieved in conventional optomechanical system (i.e., $E_{N}=0.35$Vitali et al.
+In dissipative coupling case ($g_{\omega}=0$), we observe that the optimal entanglement can be significantly stronger than in the coherent coupling case, even surpassing the entanglement achieved in conventional optomechanical systems (i.e., $E_{N}=0.35$Vitali et al.
 (2007).
@@
 4(c).
-We observe a sharp decrease in entanglement as the system temperature increases from 0 K to 5 K.
+We observe a sharp decrease in entanglement as the system temperature increases from 00 K to $5$ K.
 However, beyond this point, the entanglement becomes more tolerant to further temperature increases, suggesting that the dissipative coupling becomes more robust to thermal noise at higher temperatures.
-Remarkably, entanglement persists even up to 25 K, This indicates that strong, noise-resistant entanglement can be generated through dissipative optomechanical coupling, opening new possibilities for quantum information processing and communication in noisy
+Remarkably, entanglement persists even up to $25$ K.
+This indicates that strong, noise-resistant entanglement can be generated through dissipative optomechanical coupling, opening new ways for quantum information processing and communication in dissipative systems.
 V.3 Cooperative coupling ($g_{\omega}\neq 0$ and $g_{\kappa}\neq 0$)
@@
 5(a).
-The reduction in entanglement possibly arises from the interference between the coherent and dissipative couplings.
-When the bath temperature increases to $T=4$ K, it is observed that the optomechancial entanglement is further suppressed, as depicted by the blue dashed curve in Fig.
+This degradation of entanglement results from the destructive interference between the coherent and dissipative coupling mechanisms.
+When the bath temperature increases to $T=4$ K, the entanglement is further reduced, as depicted by the blue dashed curve in Fig.
 5(b).
-Moreover, the larger $g_{\kappa}$ is not benificial to entanglement generation.
-This can also be revealed by Fig.
-5(c), where $g_{\kappa}/g_{\omega}=0.01$ (the red curve) and $g_{\kappa}/g_{\omega}=0.1$ (the blue curve) are taken .
-For $g_{\kappa}/g_{\omega}=0.1$, the survival temperature of the entanglement is about $5$ K.
-But for $g_{\kappa}/g_{\omega}=0.01$, the survival temperature can be up to $8$ K.
+Furthermore, a larger $g_{\kappa}$ is found to be detrimental to the generation of entanglement.
+This is further illustrated in Fig.
+5(c), where $g_{\kappa}/g_{\omega}=0.01$ (the red curve) and $g_{\kappa}/g_{\omega}=0.1$ (the blue curve) are considered.
+For $g_{\kappa}/g_{\omega}=0.1$, the survival temperature of the entanglement is about $5$ K, whereas for $g_{\kappa}/g_{\omega}=0.01$, the survival temperature can reach up to $8$ K.
 VI Conclusion
-In a summary, we investigate quantum entanglement in a dissipative optomechanical system, realized by a fixed mirror and a Michelson–Sagnac interferometer with a membrane.
-This system allows the dissipative and coherent couplings to be independently tuned on and off by adjusting the membrane position or the reflectivity (transmissivity) of the beam splitter, enabling us to explore quantum entanglement in three distinct scenarios: coherent coupling case, dissipative case, and cooperative case.
-Our results reveals that, in the coherent coupling case, the steady-state mechanical displacement exhibits a nonlinear dependence on the driving power, whereas in the dissipative coupling case, the displacement shows a linear dependence.
+Before concluding, we find it worthwhile to briefly discuss the conditions under which the coherent and dissipative coupling strengths, $g_{\omega}$ and $g_{\kappa}$, can become comparable or even equal, particularly in the context of implementing either purely coherent or purely dissipative coupling mechanisms.
+In our setup, based on the Michelson-Sagnac interferometer configuration Xuereb et al.
+(2011); Tarabrin et al.
+(2013); Sawadsky et al.
+(2015); Friedrich et al.
+(2011), the single-photon coherent and dissipative coupling strengths can be independently tuned by adjusting the reflectivity of the beam splitter and the position of the membrane.
+This tunability allows access to both purely dispersive coupling ($g_{\kappa}=0$) and purely dissipative coupling ($g_{\omega}=0$).
+However, we emphasize that achieving equal values of $g_{\omega}$ and $g_{\kappa}$ under their respective coupling regimes involves different levels of experimental complexity.
+For instance, realizing a purely dissipative interaction typically demands more precise interferometric alignment and control compared to the conventional dispersive case.
+As a result, achieving equal values of $g_{\omega}$ in a purely coherent optomechanical system and $g_{\kappa}$ in a purely dissipative optomechanical system is generally challenging.
+Fortunately, under strong driving conditions, the optomechanical system can be linearized, resulting in effective coupling strengths that are proportional to the amplitude of the driving field [see Eq.
+(IV)].
+This enables the use of two independent MSI setups, one exhibiting purely dispersive coupling and the other purely dissipative coupling, each driven by an appropriately chosen strong optical field.
+Although the intrinsic coupling strengths $g_{\omega}$ and $g_{\kappa}$ may differ, by tailoring the driving field amplitudes, one can achieve comparable or even equal effective coupling strengths, $G_{\omega}$ and $G_{\kappa}$.
+This approach ensures a fair comparison in the investigation of quantum entanglement, as presented in Figs.
+3 and 4, despite the inherent differences in the physical realization of the coupling mechanisms.
+In summary, we investigate quantum entanglement in a dissipative optomechanical system, realized by a fixed mirror and a Michelson-Sagnac interferometer with a membrane.
+This system allows the dissipative and coherent couplings to be independently tuned on and off by adjusting the membrane position or the reflectivity (transmissivity) of the beam splitter, enabling us to explore quantum entanglement in three distinct scenarios: coherent coupling, dissipative coupling, and cooperative coupling.
+Our results reveal that, in the coherent coupling case, the steady-state mechanical displacement exhibits a nonlinear dependence on the driving power, while in the dissipative coupling case, the displacement shows a linear response.
 Furthermore, entanglement generated through dissipative coupling is not only significantly stronger but also more noise-tolerant compared to that achieved via coherent coupling.
-When both coupling mechanisms coexist, quantum entanglement is reduced, likely due to the interference between the two coupling interactions.
-These findings suggest that dissipative coupling may be a promising mechanism to engineer strong and noise-tolerant quantum entanglement.
+However, when both coupling mechanisms coexist, quantum entanglement is suppressed due to the destructive interference between the coherent and dissipative interactions.
+These findings suggest that dissipative coupling offers a promising mechanism to engineer strong and noise-tolerant quantum entanglement.
 This work is supported by the National Natural Science Foundation of China (Grant No.